### v1

# A Survey on Integrated Sensing, Communication, and Computation

Abstract
The forthcoming generation of wireless technology, 6G, promises a revolutionary leap beyond traditional data-centric services. It aims to usher in an era of ubiquitous intelligent services, where everything is interconnected and intelligent. This vision requires the seamless integration of three fundamental modules: Sensing for information acquisition, communication for information sharing, and computation for information processing and decision-making. These modules are intricately linked, especially in complex tasks such as edge learning and inference. However, the performance of these modules is interdependent, creating a resource competition for time, energy, and bandwidth. Existing techniques like integrated communication and computation (ICC), integrated sensing and computation (ISC), and integrated sensing and communication (ISAC) have made partial strides in addressing this challenge, but they fall short of meeting the extreme performance requirements. To overcome these limitations, it is essential to develop new techniques that comprehensively integrate sensing, communication, and computation. This integrated approach, known as Integrated Sensing, Communication, and Computation (ISCC), offers a systematic perspective for enhancing task performance. This paper begins with a comprehensive survey of historic and related techniques such as ICC, ISC, and ISAC, highlighting their strengths and limitations. It then explores the state-of-the-art signal designs for ISCC, along with network resource management strategies specifically tailored for ISCC. Furthermore, this paper discusses the exciting research opportunities that lie ahead for implementing ISCC in future advanced networks. By embracing ISCC, we can unlock the full potential of intelligent connectivity, paving the way for groundbreaking applications and services.
Index Terms:
Integrated sensing and communication, integrated sensing and computation, integrated communication and computation, integrated sensing-communication-computation, signal design, network resource management, and task-oriented communications.
I Introduction
Across the boundaries of fifth-generation wireless technology (5G), the next generation of wireless communications, 6G, is expected to go far beyond conventional communications of connected people and things to connected intelligence or intelligence of everything [1, 2, 3, 4]. As announced by IMT-2030, three additional usage scenarios of integrated artificial intelligence (AI) and communications, integrated sensing and communication (ISAC), and ubiquitous connectivity, will be fostered in 6G, as well as three enhanced usage scenarios, i.e., immersive communication, massive communication, and hyper reliable and low-latency communication, which evolve from the scenarios of enhanced mobile broadband (eMBB), massive machine-type communication (mMTC), and ultra-reliable low-latency communication (URLLC) in 5G [5]. To facilitate the implementation of these scenarios, higher key performance indicators (KPIs) are designed in 6G [4, 6]. For example, the user-experienced data rate should achieve 1 Gbit/s and the connection density should be $10^{7}$ devices/km2. Fig. 1 depicts the evolution of KPIs and usage scenarios. To achieve these KPIs, novel theories and techniques such as electromagnetic information theory [7], semantic communications [8], edge AI [9], and near-field communications [10], are developed. Moreover, the realization of connected intelligence demands the implementation of massive intelligent applications at the network edge such as autonomous vehicles, smart cities, remote health care, and industrial automation [9]. These applications call for the fusion of physical, biological, and cyber worlds, which naturally involves three modules, including sensing the environment for obtaining information, communication between different entities for information sharing, and computation on devices and servers for processing the information and making intelligent decisions [11]. This gives rise to a brand-new research area called integrated sensing, communication, and computation (ISCC). In the sequel, the related partial integration technologies are first introduced, followed by the discussion of the motivations and challenges of ISCC.
Figure 1: Evolution of KPIs and usage scenarios of 6G.
I-A Technologies of Partial Integration
I-A1 Integrated Communication and Computation
Integrated communication and computation (ICC) refers to the joint design of communication and computation processes in tasks by allowing resource allocation between them to improve the system performance under a unified goal, such as energy/latency minimization and computation accuracy maximization. Since a dozen years ago, researchers in the wireless communication field have made efforts to develop ICC techniques at the network edge. Initially, mobile edge computing (MEC) is proposed, which aims at offloading partial or full on-device tasks to edge servers for enhancing the system performance e.g., energy efficiency and latency, via balancing the on-device communication and computation loads [12]. Building on MEC, edge AI has attracted considerable attention, which aims at conducting distributed model training and cooperative inference at the network edge to exploit the benefits of fast mobile data access, data privacy preservation, and so on [9]. With the rapid advancement of microchip technology (the number of computer chips doubles every 18 to 24 months according to Moore’s Law), massive computing powers are deployed on cloud, edge, and devices in wireless networks. Accordingly, a new ICC paradigm called computing power networks is proposed for coordinating and scheduling the ubiquitous computing resources to accomplish computationally intensive tasks, such as distributed training of large models [13, 14]. Particularly, since the deployment of massive low-earth orbit (LEO) satellites is regarded as the solution for providing infrastructure-free areas with communication services [15], space computing power networks will be a promising research topic. Apart from the ICC schemes in the application layer, a physical layer technique called over-the-air computation (AirComp) has emerged in recent years [16]. It allows all transmitters to send their signals over the same wireless resource block and directly calculates a functional value of all signals at the receiver using the waveform superposition property, instead of decoding each transmitter’s signal. As a result, the communication efficiency is significantly enhanced. Due to its merits, AirComp is widely adopted as an air interface technique for wireless sensory data aggregation [17, 18, 19, 20], federated edge learning [21, 22, 23, 24], and edge AI inference [25, 26, 27].
I-A2 Integrated Sensing and Computation
Integrated sensing and computation (ISC) refers to the joint design of sensing transceivers and the subsequent sensing echo signal processing algorithms. The development of ISC techniques is significant since the realization of connected intelligence calls for deploying massive sensors in the network to sense the environment [28]. In this paper, we categorize ISC technologies into three types. The first is wireless (radar) sensing, which was originally developed in the military field for detecting aircraft before World War II. By carefully designing a transmit waveform, e.g., frequency modulation continuous wave (FMCW), and processing the corresponding echo signals with the algorithms of sampling, fast Fourier transform (FFT), correlation calculation, filtering, etc., the distance and velocity of the target can be detected [29]. Other waveforms such as orthogonal frequency division multiplexing (OFDM) waves, novel techniques like multiple antennas, and new frequency bands such as millimeter wave bands, can be used for enhancing the detection performance [29]. In addition, the utilization of wireless sensory data to train AI models for supporting various intelligent tasks such as human motion recognition and natural scene analysis, has become a research trend [29]. However, a single modality of wireless sensing is not able to obtain sufficient information for complicated intelligent tasks [30]. To this end, multi-modal sensing (including radar, camera, microphone, etc.) technology is developed. Benefiting from the rapid development of AI, machine learning algorithms are generally used for multi-modal sensory data processing, including representation, translation, alignment, and fusion [31, 32]. Moreover, mobile devices equipped with several types of sensors such as accelerometers, GPS, cameras, and microphones, have experienced exponential growth in the last several decades. This leads to a new paradigm called mobile crowdsensing, which collects massive and multi-domain sensory data from mobile devices for providing intelligent and people-centric services [33].
I-A3 Integrated Sensing and Communication
ISAC jointly designs the sensing and communication modules in the system, which saves hardware space and improves resource utilization by sharing hardware (e.g., transceivers) and network resources (e.g., time, spectrum, and energy), as well as making sensing and communication serve as each other’s supporting techniques to enhance their respective performance, so-called sensing for communication and communication for sensing. Recently, ISAC has attracted extensive research focuses and is regarded as one usage scenario of 6G by IMT-2030 [5]. There are three ISAC paradigms, corresponding to the three ISC technologies. The first is integrated wireless sensing and communication [34, 35]. By equipping dual-functional radar communication (DFRC) systems on devices, integrated wireless sensing and communication enjoy the advantages of saving hardware space and sharing spectrum resources. One main challenge lies in dealing with co-channel interference. Particularly, Wi-Fi sensing that utilizes Wi-Fi signals for detection, recognition, and estimation [36], has been a hot topic in this area, as the Wi-Fi equipment is widely deployed and no extra hardware is needed [37]. However, complete and fine-grained characteristics of both sensing target and wireless channels cannot be obtained from only the uni-modal electromagnetic wave. To this end, the second paradigm, i.e., integrated multi-modal sensing and communication, is proposed [38]. One of its main functions is sensing for communication, where multi-modal sensory data is used to assist channel state information (CSI) acquisition. The other is communication for sensing, where communication plays a key role in multi-agent multi-modal sensing systems by enabling data fusion and coordination. For the third paradigm, i.e., integrated mobile crowdsensing and communication, current research focuses on utilizing existing communication techniques for collecting massively distributed sensory data. Thus, it remains a potential area for future investigation.
I-B Motivations and Challenges of ISCC
I-B1 Motivations
Future wireless networks are expected to complete complicated tasks to provide intelligent services such as smart cities, auto-driving, virtual/augmented reality, and smart manufacturing [1, 2]. These tasks generally involve three modules, i.e., information perception (sensing), information transmission (communication), and information processing for intelligent decision-making (computation) [11, 39, 40]. Existing techniques, i.e., ISC, ICC, and ISAC as mentioned before, make efforts to support a part of intelligent tasks with loose performance requirements. However, they fail to achieve the extreme performance demands of real-time intelligent tasks, like ultra-high decision accuracy and ultra-low latency. Specifically, They have the following three shortages.
  • Low Resource Utilization: In existing designs, network resources are shared and coordinated between only two of the sensing, communication, and computation modules, leading to low resource utilization.
  • Mismatch between Goals of the Overall Task and Each Module: In existing designs, different modules are designed under separate design goals, e.g., sensing for obtaining high-quality sensory data, communication for throughput maximization, and computation for low energy cost and low latency. However, enhancing these goals is not necessarily beneficial or sometimes even harmful to the overall goal of the task [41]. For example, the transmission of unimportant data, like gradient vectors with zero values in federated learning tasks [42] and uninformative samples in active learning tasks [43, 44], consumes system resources (e.g., time, bandwidth, and energy) but makes little contribution to the task.
  • Ignoring the Tight Coupling of Different Modules: The sensing, communication, and computation modules are highly coupled in many tasks. For example, an online federated learning task involves sensing for data sample acquisition, on-device computation for local updating, and communication for local model uploading. The three modules compete for network resources (e.g., time and energy) but the convergence performance depends on the number of and the signal-to-noise ratio (SNR) of sensory data samples, the on-device computation speed, and the communication capacity in each round. Existing techniques ignore this tight coupling and design optimization algorithms from a partial perspective, leading to performance degradation.
To overcome the above shortages, ISCC is demanded.
I-B2 Challenges
Although ISCC design can achieve a higher performance ceiling, it encounters new challenges, as elaborated in the following.
  • Inference Management: In the case of the coexistence of multiple waveforms (e.g., sensing, communication, and AirComp signals), interference should be carefully managed to guarantee the quality of each module.
  • High Signal Design Complexity: In the case of reusing one waveform to perform dual functional of sensing and communication (see, e.g., [45]) or triple functional of sensing, communication, and computation (see, e.g., [46]), low-complexity signal processing techniques are demanded.
  • Complicated Resource Management: The radio resource management (RRM) from a systematic view to jointly coordinate the sensing, communication, and computation modules is much more complicated, due to the tight coupling among them and the enlarged problem size [47].
  • Lack of Unified Design Criteria: Different from the existing communication systems with a common design goal (e.g., channel capacity and SNR), there is a lack of a unified design criterion for intelligent tasks with customized goals. For example, federated learning tasks aim to achieve several goals of high testing accuracy, minimum convergence latency, data privacy, least energy consumption on devices, etc. (see, e.g., [48, 49]). This task-oriented property leads to a new challenge of finding suitable mathematical expressions to model the ISCC problem.
To this end, this paper will present state-of-the-art ISCC techniques to address the above challenges.
I-C Organization
The rest of the paper is organized as follows. We first conduct a comprehensive survey in terms of the partial integration technologies, i.e., ICC, ISC, and ISAC in Sections II, III, IV, respectively. Then, signal designs and RRM techniques for ISCC are discussed in Sections V and VI respectively. In Section VII, the research opportunities of implementing ISCC in future advanced networks, i.e., digital-twin-enabled networks, computing power networks, and space-air-ground integrated networks (SAGINs), are introduced. Finally, Section VIII concludes the paper. The detailed contents are listed below and are shown in Table I.
TABLE I: Contents of ISCC Survey
Sections Topics Details
Section II: Integrated Communication and Computation Mobile Edge Computing Definition: Offload on-device computation load to edge servers.
$\bullet$ Joint resource allocation between communication and computation (e.g., [50]),
$\bullet$ Interdependent task scheduling to ensure task efficiency and effectiveness [51],
$\bullet$ Cooperative computing to exploit on-device computing resources (e.g., [52]).
Over-the-Air Computation Definition: Calculate target functions of multiple simultaneously transmitted signals using waveform superposition property.
$\bullet$ Efficient data fusion (e.g., [19]),
$\bullet$ Distributed Consensus acceleration (e.g., [53]),
$\bullet$ Model aggregation in federated learning (e.g., [54]).
Section III: Integrated Sensing and Computation Wireless (Radar) Sensing Definition: Utilize wireless echo signals for ranging, detection, and recognition.
$\bullet$ Radar waveform design and signal processing (e.g., [55]),
$\bullet$ Radar architectures (mono-/bi-/multi-static [56, 57], phased array [58], and MIMO[59]),
$\bullet$ Intelligent radar sensing systems and applications (e.g., AI for radar type recognition [60], radiation characteristic identification [61], and natural scene analysis [62]),
Multi-Modal Sensing Definition: Complete computation tasks using multi-modal sensory data.
$\bullet$ Key issues: Representation [31], translation [63], alignment [64], fusion [65], co-learning [66],
$\bullet$ Applications (e.g., speech recognition [67], emotion recognition [68], video summarization [69]).
Mobile Crowdsensing Definition: Leverage sensory data collected from built-in sensors of ubiquitous mobile devices to empower intelligent and people-centric services.
$\bullet$ Advantages: abundant data and high coverage [70], low cost [71], scalability [72], fast data collection and response [73], and empower innovative and complex tasks [74],
$\bullet$ Key techniques: Data storage (centralized, distributed, structured, and unstructured) [33], signal processing [75], and task-level management [76],
$\bullet$ Incentive mechanisms: Money [77], intrinsic reasons (e.g., interest and personal fun) [78],
$\bullet$ Applications (e.g., emergency management and prevention [79], environmental monitoring & waste management [80], and E-commerce [81])
Section IV: Integrated Sensing and Communication Integrated Wireless Sensing and Communication Definition: Enable hardware and resources sharing between wireless sensing and communication.
$\bullet$ General integrated wireless sensing and communication: Radar communication coexistence (interference management via channel state information estimation) [82] and dual-function radar communication (a transceiver enabled with simultaneous sensing and communication) [83],
$\bullet$ Wi-Fi sensing: Exploit Wi-Fi signals for detection, recognition, and estimation [36].
Integrated Multi-Modal Sensing and Communication Definition: Joint design of multi-modal sensing and communication.
$\bullet$ Challenges and techniques: Dataset construction [38], CSI estimation for communication enhancement [84], and multi-modal sensing ability enhancement [38],
$\bullet$ Haptic Communications: Exchange haptic information in networked teleoperation systems for supporting remote immersive applications [85].
Integrated Mobile Crowdsensing and Communication Definition: Joint design of mobile crowdsensing and communication.
$\bullet$ Edge-network-level techniques: Collect sensory data in edge networks [33],
$\bullet$ Core-network-level techniques: Remote data transmission by using a high-speed fiber-optic architecture [33].
Section V: Signal Design for ISCC Single-Functional Signal Design Definition: Design independent signals dedicated for sensing, communication, and computation functions.
$\bullet$ Interference management among different signals [86].
Dual-Functional Signal Design Definition: Design one signal for sensing and another for ICC.
$\bullet$ Coexistence of sensing and AirComp [45].
Triple-Functional Signal Design Definition: Design one signal for ISCC.
$\bullet$ ISCC over-the-air [46].
Section VI: Network Resource Management for ISCC Joint RRM for Task Coexistence Definition: Sensing, communication, and computation tasks coexist and compete for network resources to achieve their own targets.
$\bullet$ Coexistence of sensing, communication, and computation tasks,[87],
$\bullet$ Coexistence of sensing and MEC tasks [88],
$\bullet$ Coexistence of ISC and MEC tasks [87]
Task-Oriented ISCC Resource Allocation Definition: Allocate network resources among the sensing, communication, and computation modules of a task to achieve its customized goal.
$\bullet$ ISCC for vehicular networks [89], machine communications [90], mobile crowdsensing [75], sensor networks [91], federated learning [48], edge AI inference [26].
Section VII: ISCC in Future Advance Networks ISCC for Digital Twin Enabled Wireless Networks $\bullet$ Hybrid AI training based digital twin network construction,
$\bullet$ ISCC based digital twin network application.
Computing Power Networks Supported ISCC $\bullet$ Computing power sharing for various ISCC tasks,
$\bullet$ Directed graph-based joint scheduling of sensing, communication, and computation.
Space-Air-Ground Integrated Networks Supported ISCC $\bullet$ Terrestrial sensing based ISCC,
$\bullet$ Remote sensing based ISCC.
  • Section II conducts a detailed survey of two ICC technologies, i.e., MEC and AirComp.
  • Three kinds of ISC technologies, i.e., wireless (radar) sensing, multi-modal sensing, and mobile crowdsensing, are discussed in Section III.
  • In Section IV, a comprehensive investigation of ISAC is provided, including integrated wireless/multi-modal/crowd-sensing and communication techniques.
  • Section V introduces three types of signal designs for ISCC, i.e., single-/dual-/triple-functional signal designs. Particularly, triple-functional signal design refers to using one signal to perform three functions of sensing, communication, and computation.
  • In Section VI, the network resource management paradigms for ISCC are discussed. One is joint RRM for task coexistence, which refers to resource allocation among different tasks of sensing, communication, and computation to achieve multiple goals simultaneously. The other is task-oriented ISCC resource allocation, which refers to allocating network resources among the sensing, communication, and computation modules of a task to achieve its customized goal.
  • The implementation of ISCC in future advanced networks is investigated in Section VII, including ISCC for digital twin enabled wireless networks, computing power networks supported ISCC, SAGINs supported ISCC.
  • Section VIII concludes the paper.
II Integrated Communication and Computation
II-A Overview
Emerging intelligent services over wireless networks often encompass both communication and computation processes. In conventional wireless networks, the communication and computation processes are separately designed. For instance, service tasks are typically executed at a central server with high computing and storage capabilities, where the central server first collects the data from geographically dispersed devices for task-related computation and then feeds back the computation results to corresponding devices for fulfilling the service requirements. With the proliferation of wirelessly connected devices, the communication burden of bandwidth-limited wireless networks significantly increases and the computation burden of the central server is also augmented due to massive volumes of data to be processed, making it inefficient to handle diverse service requirements in a centralized manner. Such a separate design ignores the inherent connection between the transmitted data and subsequent service tasks, leading to inefficient resource utilization and in turn degrading the system performance.
By treating the communication and computation processes as a unified module, ICC emerges as a promising technique to enhance the system performance, where the allocation of communication and computation resources are jointly optimized to achieve a common objective, e.g., energy/delay minimization and computation accuracy maximization, thereby making full use of limited resources to improve the system performance. In this section, we provide an overview of the existing studies on ICC, focusing on the efficient design of joint communication and computation, which is illustrated by presenting two promising examples, e.g., MEC and AirComp, as shown in Fig. 2 and listed below.
  • Mobile Edge Computing: MEC intends to move the data transmission and processing operations to the network edge by exploiting the computing power and storage capabilities of edge servers and distributed devices, while the communication resources can be utilized to exchange for the computation resources via task offloading, thereby enabling ICC with joint communication and computation resource allocation.
  • Over-the-Air Computation: AirComp aims at efficiently computing a target function by exploiting the co-channel interference, where the simultaneously transmitted signals can be naturally added over the air to realize a function computation during the communication process, thereby achieving ICC by exploiting the waveform superposition of concurrently transmitted signals.
Figure 2: Illustration of two typical examples of ICC, i.e., MEC and AirComp.
II-B Mobile Edge Computing
In the past few years, computation tasks over wireless networks have been typically accomplished by resorting to cloud computing, which requires the large amount of data generated by devices to be collected and transmitted to a centralized server with high storage and computing capabilities, so as to perform data analysis and computation for meeting various requirements of different devices. However, with the rapid development of Internet of Things (IoT), the number of connected devices dramatically increases, leading to a severe communication burden of the wireless network and high computing pressure of the central server. Therefore, due to limited radio resources, cloud computing is no longer suitable for latency-sensitive services with massive connectivities in future wireless networks. To address the above issues, MEC is emerging as a new computing paradigm thanks to the advancement of microchip computing power, which enables the edge server and mobile devices to execute computing tasks with their own computing and storage capabilities at the network edge [12].
A typical MEC system consists of edge servers and mobile devices, where edge servers are generally co-located with access points (APs) and deployed in proximity to mobile devices. Herein, the mobile devices are allowed to offload the computational tasks to edge servers for executing the tasks with high computing pressure and low latency requirements. To unleash the potential of MEC, numerous studies have explored task offloading design in MEC systems with single/multiple edge server(s) and multiple users. Specifically, the system design of MEC lies in the following aspects.
  • Joint Resource Allocation: Even though the edge servers are capable of providing strong computing power to fulfill various service requirements for different devices, they still confront high computation and communication burdens when there are massive tasks from multiple devices that need to be accomplished due to the limited radio and computing resources of the MEC system. To overcome the above challenge, joint communication and computation resource allocation is required to enhance the system performance in terms of different objectives, e.g., execution delay and energy consumption. For instance, the authors in [50] propose a task offloading framework for MEC in software-defined ultra-dense networks, where the task placement and resource allocation are jointly optimized to minimize the average task duration under the given battery energy capacity of each device. Besides, the authors in [92] develop a SAGIN-assisted MEC for computing offloading in remote areas lacking conventional edge/cloud infrastructure, where a learning-based joint resource allocation and task scheduling approach is proposed to reduce the execution delay. To mitigate the overhead for collecting the global information to enable centralized system design, the authors in [93] propose an efficient decentralized algorithm for the task offloading in a dense wireless network with selfish mobile devices. By using game-theoretical analysis, they prove the existence of pure strategy Nash equilibria and the computation costs, i.e., the weighted sum of delay and energy consumption, of different devices can be balanced.
  • Task Scheduling: In real-world applications, tasks are not only delay-sensitive and computation-intensive but also dependent on each other, which demands sequential task processing. During such a procedure, the output of one task serves as the input for the next one, which is required to be continued until all associated tasks are completed, thereby finishing the entire task. Therefore, the above issue makes it critical to schedule the offloading of such interdependent tasks in MEC systems, so as to ensure efficient and effective application performance. To this end, the authors in [51] investigate the task offloading scheme for multiple interdependent computation tasks in MEC systems, where the interdependent tasks are modeled as a directed acyclic graph. They formulate a mixed-integer nonlinear programming problem, and propose the highest response ratio and optimization method to implement joint scheduling and offloading design, which effectively reduces the average completion time in MEC systems with interdependent tasks. Furthermore, instead of assuming all requested tasks can be completed, the authors in [94] explore the task scheduling by taking into account the impact of interdependency among different tasks on the system reliability, where a multi-priority task sequencing algorithm and a deep deterministic policy gradient-based learning algorithm are developed to minimize the system deadline violation ratio.
  • Cooperative Computing: The mobile devices in MEC systems generally offload their computation-intensive tasks to edge servers for low-latency task execution while reducing communication overhead. Nevertheless, edge servers may become unavailable due to factors such as limited coverage area and exhausted computing resources, resulting in new tasks offloaded by devices not being processed instantly or failing directly. To address the above challenge, multi-user cooperative computing emerges as a promising solution to enhance the reliability of MEC systems by alleviating edge server overload and exploiting computing resources distributed among mobile devices. For instance, the authors in [52] propose a device-to-device (D2D)-assisted co-offloading framework for industrial IoT, and an efficient learning-based algorithm is developed to minimize the weighted sum of task delay and migration cost without the need of complete offloading information. To minimize the average delay under time-varying idle resources and channel conditions, the authors in [95] develop an online joint optimization for the task partitioning and parallel scheduling in a D2D-enabled MEC system, which achieves near-optimal delay performance with low computational complexity while ensuring fairness across various system states. In addition, the authors in [96] introduce a distributed multi-hop task offloading decision model for Internet of Vehicles (IoV) networks, where the peripheral vehicles with idle computing resources can be exploited to improve the task execution efficiency in dynamic vehicular environments.
II-C Over-the-Air Computation
In traditional communication protocols, the signals among different devices over the same channel are deemed as co-channel interference, thus each of them demands individual and accurate reception at the receiver before the subsequent processing. In the face of emerging intelligent services, achieving functional computation of gathered data becomes the key to implementing the corresponding applications, e.g., environmental monitoring and autonomous driving, while the accuracy and timeliness of such computation determine the quality of service. In this sense, the conventional compute-after-communicate strategy may become inefficient in a large-scale network with scarce resources due to the separation of communication and computation. With this in mind, AirComp has attracted much attention to pursuing the fusion of communication and computation by harnessing the co-channel interference [97, 98, 99].
Seeking to integrate communication and computation, AirComp regards the co-channel waveform superposition as a summation operation for synchronized concurrently transmitted signals, enabling compute-when-communicate manner that allows the target function to be computed during the transmission process without need for one-by-one decoding at the receiver. As all devices are required to occupy the same channel to enforce the computation for the target function, the number of radio resources utilized by AirComp is unrelated to the number of participating devices, thereby yielding a much higher spectral efficiency than conventional orthogonal multiple-access (OMA) approaches that allocate separated channels to each device. Meanwhile, compared with the non-orthogonal multiple-access (NOMA) scheme counterparts, AirComp avoids the complex design for interference cancellation and saves the time for recovering each signal in terms of the functional computation at the cost of acceptable computation errors.
By implementing appropriate pre- and post-processing on transmit and receive signals, respectively, AirComp enables the receiver to compute a class of functions, namely nomographic function [100], which is defined as $f(x_{1},x_{2},\dots,x_{K})=\psi\left(\sum_{k=1}^{K}\varphi_{k}(x_{k})\right)$, where $\psi(\cdot)$ denotes the post-processing function at the receiver and $\varphi_{k}(\cdot)$ denotes the pre-processing function for data $x_{k}\in\mathbb{R}$ at device $k$, $\forall\,k\in\{1,2,\dots,K\}$. Following the expression given in the above, AirComp can be decomposed into three stages [16]: 1) transmit data pre-processing at each device; 2) summation of pre-processed data through the waveform superposition property of multiple-access channels; and 3) receive data post-processing at the access point (AP). Note that, although the signal superposition can only realize the operation of summation, AirComp is capable of computing non-linear functions [101], e.g., geometric mean and Euclidean norm, and even computing any continuous function by assembling multiple nomographic computation results [100].
During the wireless propagation, the synchronized signals are performed with a weighted summation with respect to the channel coefficients, which may lead to a non-desired computation result. Accordingly, the transceiver design is necessary for AirComp to induce an expected signal superposition over non-uniform fading channels. Recently, many efforts have been made to induce signal alignment under different objectives for AirComp.
  • Data Fusion: Data fusion is the fundamental objective of AirComp and much of the underlying research has been conducted from this aspect [102]. For instance, the authors in [19] reveal the optimal transceiver design for AirComp under peak power constraints to minimize the distortion of data fusion in single-input single-output (SISO) networks. Meanwhile, the optimal design under sum-power constraint is obtained in [103] with closed-form solutions. By exploiting multi-antenna techniques, the authors in [16, 104, 105] develop beamforming design for AirComp and validate the performance improvement that can be achieved via diversity gain and multiplexing gain of multi-antenna arrays to reduce fusion error [104] and realize multi-modal fusion [16, 105], respectively.
  • Distributed Consensus: Distributed consensus expects all devices in the network to reach a consensus on a certain message without the assistance of a central node, which can be accelerated by AirComp with its efficient data aggregation. For instance, the authors in [53] show that multiple simultaneous average operations achieved by AirComp can accelerate the convergence rate of gossip-based consensus. In addition, the authors in [106] propose beamforming design for decentralized AirComp to accelerate the distributed consensus.
  • Model aggregation: Federated learning (FL) has attracted much attention in machine learning areas due to its distributed computing structure and privacy-preservation property [54]. Although transmitting model parameters reduce the communication overhead by a significant amount compared to transmitting raw data, scarce communication resources still limit the communication efficiency in wireless FL systems, which motivates the emergence of AirComp-based FL [21, 22, 107, 108]. The authors in [21] demonstrate that AirComp-based FL achieves much lower communication latency than the OFDMA-based scheme. To ensure accurate model aggregation, the authors in [22, 21, 109] develop device scheduling methods. Moreover, compression and sparsification approaches are leveraged in [110] to further reduce the communication overhead in resource-constrained wireless networks.
III Integrated Sensing and Computation
III-A Overview
In this section, we provide a comprehensive survey about the technologies for both sensing signals acquisition and the corresponding processing methods for completing sensing tasks, giving its name ISC. As shown in Fig. 3, there are three ISC technologies.
Figure 3: Sensing and the corresponding signal processing techniques.
The first one is wireless sensing (also called radar sensing) technology, which senses a target by receiving and processing the echo signal of a carefully designed waveform. Yet, information obtained from a single modality of sensing signals may not be sufficient for completing complicated tasks [111]. It is mathematically shown in [30] that collecting multi-modal sensory data can provide better sensing performance. Therefore, the second ISC technology of multi-modal sensing is developed. Moreover, although multi-modal sensing technology can enjoy many benefits of rich information, high robustness, cross-modality data transformation, it requires deploying dedicated sensors, which is expensive. To empower ubiquitous sensing in a large coverage with low cost, the third ISC technology called mobile crowdsensing, is proposed [33]. It utilizes built-in sensors of ubiquitous mobile devices for collecting sensory data in various domains. In the sequel, the above three ISC technologies are discussed.
III-B Wireless Sensing Technology
Wireless sensing, which is also called radar sensing, emits modulated radio waves, receives and processes the corresponding echo signals to complete computation tasks such as object ranging and detection, and motion recognition. Its application can be traced back to the usage of radar in the early 20th century. Because of its significant military value, radar ushered tremendous development during World War II and turned to be a key technique for empowering many major and intelligent applications, including military surveillance, remote sensing in space, and auto-driving. Based on the original continuous-wave radar, new techniques in terms of wave-form design, signal processing, and system architecture are developed to enhance its abilities regarding velocity, distance, direction, and sensitivity toward motions. In this part, the development of radar system design and the corresponding signal processing are introduced.
III-B1 Radar Waveform Design and Signal Processing
Aiming at achieving higher accuracy of the ranging and detection results, one key advancement of radar techniques is the development of waveform design and the corresponding signal processing approaches [55]. In this part, we will introduce the evolution of the waveform design from the initial continuous wave to the state-of-the-art designs.
  • Traditional Continuous-wave Radar: As its name suggests, traditional continuous-wave radar continuously transmits a waveform with a stable frequency and receives an echo signal from one or more targets. To detect the Doppler shift of the received echo signal, a conjugate mixing method is utilized. To be specific, both the transmit signal and the echo signal are sampled with a time interval $T$ and then transformed into the frequency domain using fast Fourier transform (FFT). By calculating the cross-correlation of the two signals and applying a low-pass filter, the frequency difference of the two signals, i.e., the Doppler frequency shift (DFS), can be extracted. Based on the DFS, the velocity can be estimated. The resolution (or the detection error) of the velocity is proportional to the sampling rate $\dfrac{1}{T}$. However, the measurement of range information i.e., distance of the target which is decided by light speed and the round-trip delay, is not available with continuous waveform. The reason is that the round-trip delay is detected by the phase shift of the echo signal, which is ambiguous to the receiver due to the lack of phase basis. That’s to say, we cannot distinguish the phase shifts $\theta$ and $(\theta+2k\pi)$ with $k$ being any integers.
  • Pulsed Wave Radar: To obtain the range information, the pulsed radar is developed, where the transmit waveform is generated by modulating a single-frequency signal by a periodic window signal with period $T$. That’s to say, the transmitter periodically produces a brief pulse of radio signal and keeps silent to receive the echo signal of the detected object. The interval between two consecutive pulses is called the inter-pulse period (IPP). The range information is detected by processing the echo signal within one transmission period, called fast-time processing. Specifically, with the time basis of the transmit signal, the exact phase shift of the two signals can be detected using their correlation in the frequency domain. Subsequently, the round-trip delay is calculated and is further used to compute the distance. It should be emphasized that IPP should be larger than the round-trip delay to make sure the echo signal of one pulse can be correctly received. The resolution of rang information is proportional to the pulse transmission period $T$. The velocity is detected using multiple pulses, called slow-time processing. With a fixed velocity, the DFS of each signal in several consecutive pulse waveforms is a single-frequency signal. Hence, by calculating the correlation of the echo signal and the transmit signal over a long time scale (i.e., over multiple transmission periods) in the frequency domain, the DFS is detected and the corresponding velocity can be calculated. The resolution of the detected DFS or velocity is proportional to $1/T$. Note that there is a tradeoff between the resolutions of range and velocity information.
  • Frequency Modulated Continuous Wave (FMCW) Radar: The time efficiency of pulsed wave radar is relatively low due to the intermittent transmission. To this end, an advanced FMCW radar is developed. The sensing signal is decomposed of multiple repeated snapshots. In each snapshot, a continuous waveform is transmitted with the frequency modulated as a linear function of time. Although transmitting and receiving work simultaneously, no interference is caused as they have different frequencies. Similar to pulsed radar, the range information is detected via fast-time processing, and the velocity is detected by slow-time processing using $M$ snapshots, but the time efficiency of FMCW is higher compared to pulsed radar due to continuous transmission. The resolutions of range and velocity are proportional to the snapshot duration and the number of used snapshots $M$, respectively.
III-B2 Radar Architectures
In this part, five radar architectures are introduced, as shown in Fig. 4.
Figure 4: Illustration of five radar architectures.
According to the locations and numbers of transmitters and receivers, radar architectures are categorized into mono-static, bi-static, and multi-static types with each having its pros and cons [56, 112, 57, 113], which are discussed as follows.
  • Mono-static Radar: The mono-static radar refers to that the transmitter and receiver are co-located [56, 112]. As a result, this architecture enjoys a concise structure and can be deployed conveniently. More importantly, signal processing is simplified as the synchronization of transmit and receive signals is easy. However, due to the simple structure, its radar cross section (RCS) is limited, leading to a relatively poor performance. Besides, continuous waveforms are not feasible in mono-static radar as they suffer from severe self-interference.
  • Bi-static Radar: In a bi-static radar system, one transmitter and one receiver are separately deployed [57, 113]. Such systems are featured by the transmitter-target-receiver triangulation. Under this setting, no self-interference is caused at the receiver, leading to a more flexible choice of waveforms. The transmit and receive antennas can be implemented and optimized independently and the RCS-like range and angular resolutions can be improved via spatial filtering. However, the separate design brings more difficult time synchronization between the transmitter and the receiver. Additional information like locations and waveform types should be shared, leading to a higher cost for maintenance.
  • Multi-static Radar: Multi-static radar systems are the extension of bi-static radar systems [56]. They are consisted of two or more transmitters and receivers with these units separated by large distances when compared to the antenna sizes. Benefiting from this architecture, the probability of target loss is reduced since multiple observations of the target from different aspect angles are enabled. Besides, the energy from a target echo that is created by one transmitter can be utilized by multiple receivers, leading to reduced overall energy consumption for a large sensing coverage as well as less mutual interference. However, this architecture is more complicated, resulting in more expensive costs for time synchronization, information sharing, and joint optimization.
To improve radar’s sensing ability for object detection in a wider area, early radar systems scan the target via the rotation of a single antenna [114]. In this system, an antenna periodically emits a probing signal and receives the reflected signal (i.e., the same antenna is often used for both transmitting and receiving). As a result, time-division duplex should be carefully designed and waveforms like CW and FMCW cannot be applied. Benefiting from the advancement of multi-antenna techniques, new solutions are proposed for improving the detection coverage of radar, as described below.
  • Phased Array Radar: Phased array radar comprises multiple antenna (or radiating) elements arranged in a particular order on a curved or plane surface. Its transmit waveform beam can be electronically steered to points in different directions and distances via controlling the power and phases among these antennas without changing their locations [58, 115, 116]. For example, by letting two radiators transmit a signal in the same phase shift, a waveform superposition is achieved and the signal strength is enhanced in the main direction and attenuated in other directions. By carefully designing the power and phase control strategy, phased array radar enjoys the advantages of high antenna gain with great side-lobe attenuation, arbitrary directions scanning, high beam agility, multi-functional operations via simultaneous generation of multiple beams, and enhanced robustness, which allows failure of some components. However, the complicated structure that needs to simultaneously control all antennas leads to higher costs and problems like low-frequency agility, deformation of the antenna pattern during beam steering, and limited scanning range.
  • Multiple-input and Multiple-output (MIMO) Radar: In MIMO radar systems, multiple antennas are equipped at both the transmitter and the receiver [59, 117, 118]. In such a system, each transmit antenna radiates arbitrary waveforms independent of other transmit antennas. At the receiver, the echo signal can be mapped with a single transmit antenna due to the independence of waveforms. By paying the cost of higher signal processing complexity, MIMO radar can design more adaptive transmit/receive beamforming and more flexible probing signals to achieve higher detection resolution and better parameter identifiability [119, 120, 121].
III-B3 Intelligent Radar Sensing Systems and Applications
Radar sensing shows its superiority compared with other sensing techniques in many aspects, including long detection distance, high penetration property, easy to be equipped in vehicles. As a result, it is widely used in many military applications including military target detection, tracking, and recognition, and civil applications such as disaster assessment, land use and classification, and weather detection [62]. However, as the demands of empowering many intelligent services like auto-driving, human detection, and activity classification, modern radar systems encounter new requirements in terms of high accuracy, robustness, and real-time ability. This is beyond the capability of traditional radar signal processing techniques. To tackle this problem, AI, which has shown powerful abilities in many other areas like signal processing in biomedical and communication systems (see, e.g., [122, 123]), is applied for enabling intelligent radar signal processing.
One trend for intelligent radar signal processing aims at recognizing the radar radiation sources by processing the received waveforms with AI models [60], including two application scenarios. One is to distinguish different radar types, e.g., military radar or civil radar [61]. The other is to identify individual radar emitter parameters including the radiation characteristic, qualification of components, and so on [124]. The data used for AI model training includes pulse description words [125], amplitude-phase shift [126], I/Q time sequences [127], autocorrelation features [128], etc. Generally, to enhance the training performance, these raw data are pre-processed via e.g., denoising [129], missing data processing, and features extraction [127].
More literature about intelligent radar signal processing directly utilizes AI models to process the radar image, which refers to the echo wave power distribution in the delay-Doppler domain, for target detection, recognition, and natural scene analysis [62]. For example, AI models are used for canceling the interference and noise in synthetic aperture radar (SAR) images (see, e.g., [130]) and target detection or recognition (see, e.g.,[131, 132]). Particularly, adopting machine learning algorithms for processing micro-Doppler and range-Doppler images is widely used in many intelligent applications such as unmanned aerial vehicle (UAV) detection [133], human activity classification [134], and gait recognition [135].
III-C Multi-modal Sensing Technology
A sensory modality refers to the way one entity is observed or perceived, e.g., via radar, camera, microphone, etc. Compared with unimodal sensing like radar sensing, multi-modal sensing can collect more information to achieve better task performance, which is mathematically proved in [30]. Besides, the common information can be obtained by different modalities, making multi-modal sensory data more robust to noise and sensory deficiency, i.e., data of one modality can provide complementary or supplementary information to others [31]. For example, in an audio-visual speech recognition task, the visual information helps enhance the speech recognition performance when the audio signal is noisy [31]. Moreover, the success of the recent deep generative models enables creating one modality of sensory data from another [111], such as using a text to generate a high-resolution image.
To enjoy the above advantages and based on the fact that data in the real world have multiple modalities, multi-modal sensory data has been the fuel to empower AI in many applications, including scene application, action recognition, video transcription, image captioning, and so on [31, 111, 136]. On the other hand, data of different modalities is heterogeneous in the forms (e.g., voice, image, and text) and the noise levels. The relation between different modalities is complex. Machine learning algorithms are currently the most powerful and widely adopted methods for multi-modal signal processing. These give rise to a new research area namely multi-modal machine learning. Subsequently, the key issues of multi-modal machine learning are discussed, followed by the introduction of the applications of multi-modal sensing.
III-C1 Key Issues in Multi-modal Machine Learning
As stated in [31, 32], there are five key issues to be addressed in the area of multi-modal machine learning, namely representation, translation, alignment, fusion, and co-learning. They are elaborated below.
  • Representation: Representation is the information distilled from the data that intelligent systems can use [137]. Generally, representation and feature are interchangeably in AI tasks [31]. Many efforts have been put into efficiently extracting useful representations for AI tasks based on unimodal data, e.g., representation learning [137], feature detection [138], convolutional neural networks based feature extraction [139]. However, representation extraction from multi-modal sensory data faces new challenges. On one hand, data in each modality is obtained from different types of sensors with heterogeneous forms and noise levels. On the other hand, in many scenarios, data in some modalities is missing. The authors in [66] pointed out that a good multi-modal representation should satisfy the following novel requirements: 1) Similarity in representation should lead to the same similarity in the corresponding results; 2) Multi-modal representation should be readily obtained even in the absence of certain modalities; 3) Missing modalities can be generated using other observed ones. As a result, it remains a challenging problem to construct a good representation for multi-modal data.
    The current methods for multi-modal representation can be categorized into two types, i.e., joint representation and coordinated representation [31, 32]. Joint representation projects data of all modalities into the same latent representation space, which is the most adopted method [31]. As well as using the conventional neural networks, autoencoders are utilized for pre-training the unlabeled data (e.g., [67, 140, 141]) and probabilistic graphical models like deep Boltzmann machines are adopted to deal with the issue of missing modality via using their generative nature [66]. In coordinated representation, different modalities have their own representations but are coordinated via setting a constraint of similarity (e.g., Euclidean distance) or structure (e.g., partial order) [142].
  • Translation: Multi-modal translation refers to the ability of modality transfer i.e., generating some modalities of one entity with one source modality [31]. There are two kinds of translation approaches, i.e., example-based and generative approaches.
    • Example-based approach: In this approach, a dictionary is constructed using the training data samples. For each entity, the source modality and target modality of one entity are stored in the dictionary with one-to-one correspondence [31]. The simplest way called the retrieval-based model for modality translation is finding the closest entry in the dictionary and using it as the translation result [63]. Another way, called the combination-based model, takes one step further. It combines multiple entries in the dictionary with a designated way to create a better translation result [31]. For example, the authors in [143] find the most similar $k$ items of one source image and utilize their captions to generate a target descriptive sentence of the source image. Although the idea of the example-based approach is simple, it calls for expensive costs for the construction and inference of the dictionary especially when the number of items is large [31].
    • Generative approach: In this approach, generative AI models are trained for modality translation [31]. Specifically, two kinds of methods are included. One is called the grammar-based model, which first extracts high-level concepts from the source modality. Then, these concepts are integrated via a pre-defined grammar-based generation procedure for creating a target modality [144]. By using predefined templates and restricted grammars, this method can generate syntactically or logically correct target instances. However, this method prefers creating formulaic rather than creative translations, and a separate model and training dataset are required to detect each concept, leading to complex pipelines for concept detection [31]. The other method is the encoder-decoder model. It first encodes a source modality into a vectorial representation and then utilizes a decoder to generate the target modality [145, 146]. Although this method calls for large quantities of data for training, it is widely used in many generative tasks, like text, image and sound generation, image caption, and video description.
  • Alignment: Multi-modal alignment is defined as finding the relations and correspondences among several modalities of one entity [31]. With good alignment, the performance of modality translation can also be enhanced. Multi-modal alignment can be divided into two categories, i.e., explicit alignment and implicit alignment. The former directly aligns the sub-components among different modalities, e.g., aligning a movie to a script. The latter latently aligns data of different modalities during AI model training, targeting enhancing the performance of many intelligent tasks, like speech recognition, machine translation, etc. There are three main issues to be addressed for multi-modal alignment. The first is that not all elements in one modality have correspondences in another, especially in the case of missing modalities. The second is the difficulty in designing a similarity metric for different modalities. The third is the scarce datasets with explicitly annotated alignments [31]. In the sequel, the methods for explicit and implicit alignment are introduced.
    • Explicit alignment methods: For the case where there are no alignment labels, i.e., labeled correspondences between two modalities, unsupervised explicit alignment algorithms are required. One basic approach is called dynamic time warping (DTW) [64]. It measures the similarity between the time sequences of two modalities by time warping and aims at finding an optimal match between them. The alignment performance of such an approach highly depends on the similarity metric. Then, an extended version of DTW is proposed by introducing canonical correlation analysis (CCA) [147, 148]. As a result, both alignment (through DTW) and learning the mapping (through CCA) between the sequences of different modalities are enabled. The DTW approach and its extension align multi-modal data via linear transformations. To deal with the non-linear relations, a deep canonical time-warping approach is proposed in [149]. To further learn the complex alignment relations between modalities, various kinds of graphical models are also popular for multi-modal data alignment, including generative graphical models [150], factored hidden Markov models [151], dynamic Bayesian networks [152], etc. For the case where there are alignment labels, supervised explicit alignment algorithms are proposed, most of which take inspiration from the unsupervised methods mentioned above, in addition to utilizing the alignment labels to train a better similarity measure (e.g., [153, 154, 155]).
    • Implicit alignment methods: As mentioned, implicit alignment methods try to latently align multi-modal data for completing a machine learning task instead of directly aligning the instances of multiple modalities. The early methods adopted graphical models [156], where the mapping between modalities should be manually constructed. Other methods directly use the neural networks for implicit alignment. Specifically, various kinds of different attention models have been proposed for better performance in question answering and better model interpretability [157], including hierarchical [158], stacked [159], and episodic memory attention [160].
  • Fusion: Multi-modal fusion refers to integrating information from multiple modalities of the same phenomenon for completing a prediction task, such as classification and regression [31]. As mentioned, using multi-modal information for task completion leads to more robust performance due to their complementary property. However, there are two challenges [31]. On one hand, the signals of different modalities may not be temporally synchronized and exhibit different noise levels. On the other hand, it is difficult to build up models that can explicitly exploit the complementary properties of multiple modalities. To address the above challenges, three types of multi-modal fusion approaches are proposed, i.e., early, late, and hybrid fusion [65]. Early fusion refers to integrating the representations or features of multi-modal data and the integrated representations or features are used for completing the downstream task. It can learn the correlation and interactions of different modalities in the early stage and only a single model is trained. Late fusion is similar to ensemble learning. It fuses the decisions based on unimodal data using mechanisms like averaging [161], voting schemes [162], or a learned model [163, 164]. Late fusion uses one predictor for each modality, allowing more flexibility and is more robust to modality missing. Hybrid fusion, as its name suggests, combines the above two types to enjoy the advantages of both.
  • Co-learning: Co-learning refers to exploiting knowledge from a (resource-rich) modality to aid the modeling of another (resource-poor) modality when the latter is lack of data, polluted, or has unreliable labels [31]. To achieve high co-learning performance, several challenges should be overcome. First, the learned representation of the resource-rich modality may be task-relevant, leading to a limited performance when transferring its knowledge across different modalities in different tasks. Second, the instance numbers and labels are unbalanced. Moreover, the noise levels and the reliability of the labels are heterogeneous. According to whether the multi-modal instances of the same entity are aligned during training, the co-learning approaches are categorized into three types [31].
    • Aligned instances: In this case, two methods are adopted. One is co-training, where instances of two modalities bootstrap each other with labels for unlabelled instances [31]. This method, however, may lead to unbiased training samples and cause overfitting. The other method uses transfer learning-based multi-modal representation learning, which transfers the representation of one modality to another (e.g., [66]).
    • Unaligned instances: In this case, the training instances of different modalities are not necessary from the same entities. The first method uses transfer learning to achieve coordinated multimodal representations, e.g., using text messages to improve visual representations for image classification [165]. The second is conceptual grounding. It learns semantic meanings or concepts based on data of multiple modalities, including language, vision, and sound, by finding a common latent space of all multi-modal representations (e.g., [166]). The last one is zero-shot learning, which recognizes a concept without explicitly seeing any examples of it [31]. For example, it can use a neural network to map the text-based message with the extracted visual features.
    • Hybrid case: In this case, approaches use a pivot modality to learn coordinated multi-modal representations using e.g., bridge correlational neural network, or rely on the existence of large datasets from a similar or related task [31].
III-C2 Applications
Multi-modal sensing has empowered various kinds of applications. Via the utilization of audio and visual data, the performance of speech recognition and synthesis [67, 167], human action classification [168], multimedia event detection [169], emotion recognition [68], can be greatly enhanced compared with using unimodal information. By utilizing the modality translation techniques, many media description applications have been well developed, including image and video captioning [69, 170], visual question-answering [171], etc. By creating structured representations of multi-modal data, multimedia content indexing, and retrieval applications are enabled, including automatic shot-boundary detection [172] and video summarization [173]. Furthermore, multi-modal data can enhance the understanding, analysis, and decision-making processes in the domain of healthcare, fitness, and general wellness [174, 175]. Autonomous systems utilize various kinds of sensory data, including radar, stereoscopic visible-light cameras, light detection and ranging (LiDAR), infrared cameras, a global positioning system (GPS), and audio for enhancing the automation capability [176, 177].
III-D Mobile Crowdsensing Technology
Mobile crowdsensing, also called crowdsensing or mobile crowdsourcing, has been a powerful paradigm for leveraging built-in sensors of ubiquitous mobile devices like accelerometers, GPS, cameras, and microphones for data sensing and collection in various domains, so as to empower a large number of intelligent and people-centric services [33]. In this sequel, we discuss the advantages, key techniques, incentive mechanisms, and applications of mobile crowdsensing.
III-D1 Advantages
Compared to the traditional sensing paradigm using dedicated sensors, mobile crowdsensing demonstrates its excellence in the following aspects.
  • Abundant Data and High Coverage: The rich set of built-in sensors of massive mobile devices utilized in crowdsensing can collect abundant data and have a much larger sensing coverage [70].
  • Low Cost: Benefiting from utilizing the existing infrastructure of mobile devices, crowdsensing is more cost-effective in terms of deployment and maintenance [71].
  • Scalability: With billions of mobile devices, crowdsensing is scalable to obtain data from a diverse range of geographical locations, demographic groups, and environmental conditions, making it a promising method for addressing global issues and societal challenges [72].
  • Fast Data Collection and Response: As mobile devices can continuously acquire and transmit data, crowdsensing enjoys the benefits of real-time data collection, immediate analysis, and fast response to emergent and critical missions [73].
  • Innovative and Complex Tasks: Crowdsensing can empower a wide range of innovative and complex tasks like smart cities, social sensing, and environmental monitoring, via collecting data from multiple resources, different modalities, and various domains (see, e.g., [74, 178, 80]).
III-D2 Key Techniques
The efficient deployment of the mobile crowdsesing paradigm depends on the techniques of data collection, data storage, data transmission, signal processing, and task-level management [33]. As the data transmission techniques involve both sensing and communication, we leave its introduction in the next section and the other techniques are introduced as follows.
  • Data Collection: In mobile crowdsensing systems, data collection requires the engagement of massive user participation, where some non-professional users are included. The target of data collection techniques is acquiring reliable and trustworthy data via sensor selection. There are two frameworks. One is the domain-specific framework designed for a dedicated task such as intelligent transportation and health care [33]. As a result, the performance of the specific task can be enhanced. The main challenge of such a framework lies in the co-existence of many tasks or domains degrading the efficiency and reliability of the overall crowdsensing system. The other is a general-purpose data collection framework. As its name suggests, it is designed to support many tasks at the same time by maximizing a set of parameters or minimizing the costs [179].
  • Data Storage: The practical storage location in crowdsensing systems is task-oriented and is classified into centralized storage and distributed storage. The former stores data in a single location, typically a cloud-based database, which is applied when significant processing or data aggregation is needed, e.g., in urban monitoring and price comparison [33]. This approach requires high communication throughput for data transmission and violates the data privacy of users but can enjoy a low processing latency. Distributed storage is employed for delay-tolerant applications such as mapping air quality and fog computing, but has low communication cost and can keep users’ data privacy [180]. The storage format can be structured and unstructured [33]. In a structured format, the data has a specific identifier and is easy to access and self-explanatory, e.g., the information of individuals stored in terms of their names or identifications. In an unstructured format, there is no specific identifier, leading to difficulty in recognizing the data, e.g., the storage of videos, audio, and images.
  • Signal Processing: Signal processing in crowdsensing refers to the manipulation and analysis of the collected data. It involves three sequential processes, i.e., pre-processing, analytics, and post-processing [33]. The main target of pre-processing is filtering and denoising by removing irrelevant and redundant information, suppressing noise, aggregating data, and compressing data size to save storage volume [75]. Analytics aims to extract meaningful information from the collected data. Machine learning and data mining are common approaches for information inference, pattern recognition, or future trend prediction. For real-time tasks such as emergency situation management, traffic monitoring, and unmanned vehicles, high computational resources are needed for fast decision-making and response [33]. After analytics, post-processing can be conducted in an offline way for statistical or predictive analysis [181].
  • Task-level Management: Since there are a wide variety of crowdsensing tasks, it’s significantly desirable to assign and manage the tasks to the participating users or mobile devices. Specifically, task-level techniques include task assignment and scheduling [33].
    • Task assignment: Task assignment involves two aspects. One is task partitioning, which divides complex crowdsensing tasks into smaller and manageable sub-tasks, that can be completed by different devices [76]. For example, in [182], the overall target area is divided into multiple distinct sub-areas, with each being assigned to one device for sensing. The spatial and temporal correlation of different devices’ sensory data are leveraged. In [183], K-means algorithm is used to partition a crowdsensing task into several sub-ones, each of which is allocated to one mobile device according to their sensing abilities. The similarity of the sub-tasks is further considered in [184] during task assignment.
      The other task assignment technique is task allocation, which involves distributing tasks to users with the goal of reducing resource costs and enhancing task performance. This technique is categorized into two groups, i.e., centralized methods and distributed methods. In the former, a central coordinator assigns the tasks to users or devices, leading to an enhanced systematic performance but requiring global information (see, e.g., [182, 185]). In the latter, each device makes their own decisions during the sensing task, resulting in better data privacy preservation (see, e.g., [186]).
    • Task scheduling: Task scheduling in mobile crowdsensing refers to the way that mobile devices collect and contribute their sensory data. In proactive scheduling, mobile devices actively obtain sensory data without a pre-assigned task and can decide where and when to contribute their data. Proactive scheduling is commonly conducted for people-centric tasks (see, e.g., [187, 188]). In reactive scheduling, the sensing task and target are clearly pre-defined, which is adopted in many works. For example, the authors in [189] propose a minimum energy multi-sensor task scheduling algorithm that coordinates multiple tasks for the same user under the targets of enhancing sensory data quality and minimizing energy consumption. A first-in-first-out scheduling strategy and an arbitrary deadline (AD) scheduling strategy are respectively proposed for offline and online settings in [190]. A deep reinforcement-learning-based real-time scheduling algorithm is proposed in [191].
III-D3 Incentive Mechanisms
The crucial condition for the success of mobile crowdsensing is the ability of the network organizer to collect data from the sensors on each user’s mobile phone. This, however, violates the data privacy of mobile users and requires additional on-device power consumption for data transmission [192]. Taking the noise monitoring task as an example, continuous collection of ambient sound violates personal data privacy and reduces devices’ battery life. To encourage as many mobile users as possible to contribute their sensory data with high accuracy and reliability, incentive mechanisms become promising methods to balance the users’ additional cost and the overall performance gain of crowdsensing systems [77, 193].
Based on whether financial payment is applied, the incentive mechanisms can be categorized into monetary and non-monetary [77].
  • Monetary Incentives: Paying for sensory data is the most straightforward incentive. Monetary incentives serve as an important motivator when crowdsensing tasks don’t yield immediate benefits for users. Some research has shown that the increasing amount of payment can help the sensing task be completed faster (see, e.g., [194]). Monetary incentives are further divided into static mechanisms and dynamic mechanisms. In the former mechanisms, the reward quantity is pre-determined based on a specific criterion, and this amount remains consistent throughout the entire experiment (see, e.g., [195]). The latter mechanisms allocate a variable budget for each task, contingent on the real-time conditions of the system (see, e.g., [196]).
  • Non-Monetary Incentives: As its name suggests, non-monetary incentives have no financial payment and call for voluntary participation [197]. This severely restricts its scope of application. In some specific tasks, contributing users can benefit from the overall crowdsensing tasks, making it possible to design appropriate incentive mechanisms, including collective motives, social reward, and intrinsic motives and fun [77]. Collective incentives encourage the achievement of common goals, which is the philosophy of crowdsensing, to work together for a common good. However, once the goal is reached, non-participants can also make a profit. Hence, it’s desirable to add additional incentives to avoid “a free ride”. Social rewards are very useful when collecting sensory data from social networks. The rewards refer to the benefits that can be obtained by social networks (e.g., Facebook and TikTok) including interaction, and self-interests (see, e.g., [198, 199]). In non-monetary incentives, the inherent interest and personal fun, like gaming, are the motivation of the majority of participants [78]. This intrinsic reason can attract the target participants to contribute their own data.
III-D4 Applications
By leveraging the rich data collected via mobile crowdsensing, a wide range of applications are empowered. Several typical applications are listed below.
  • Emergency Management and Prevention Applications like accidents and natural disasters, where urgent decision-making and fast response are required [79].
  • Environmental Monitoring & Waste Management Applications such as air quality [80], noise pollution [200], and waste-recycling operations [201], where the resources, infrastructures, and phenomena, that related to people’s daily life are monitored.
  • E-commerce Applications that collect, process, and deliver prices of goods in the real physical world [81].
  • Indoor Localization Applications that provide localization and navigation services when users are in indoor environments [202].
  • Health Care and Wellbeing Applications that diagnose, treat, and prevent illness, diseases, and injuries using the collected crowdsensing data [203].
  • Intelligent Transportation Systems for efficient transportation and traffic management via monitoring traffic, vehicles, public transport, road conditions, and arrival time [204].
  • Mobile Social Networks that build social relations and meet and share data between people with similar interests using mobile devices [205].
  • Public Safety Applications for evaluating the safety exploiting the urban environments [206].
  • Unmanned Vehicles to enable interaction between mobile users and automobile vehicles [207].
  • Context Awareness that combines mobile activity recognition and data mining to detect context for benefiting data accuracy and battery maintenance [208].
IV Integrated Sensing and Communication
IV-A Overview
ISAC has been a promising technique in wireless networks for saving network resources and enhancing the performance of both sensing and communication. Different from existing literature (e.g., [29, 209, 36]) focusing on integrated wireless sensing and communication, a comprehensive survey about ISAC is provided in this work. Specifically, three ISAC paradigms will be presented which align the three ISC technologies in Section III, as shown in III and listed below.
  • Integrated Wireless Sensing and Communication: This paradigm enables hardware and radio resources sharing between wireless sensing and communication. The general integrated wireless (radar) sensing and communication technique is first elaborated, including two frameworks of radar communication coexistence (RCC) and dual-functional radar-communication (DFRC) [209]. Then, a special case of Wi-Fi sensing is discussed, since WiFi signals are widely distributed and can be used for sensing at low cost.
  • Integrated Multi-modal Sensing and Communication: Utilization of only uni-modal electromagnetic wave for ISAC cannot provide complete and fine-grained characteristics of both sensing target and wireless channels [38]. This motivates the emergence of integrated multi-modal sensing and communication. In this work, three key techniques are introduced, i.e., dataset construction, multi-modal sensing for channel estimation, and communication for multi-modal sensing enhancement.
  • Integrated Mobile Crowdsensing and Communication: The current research focus on this topic lies in using existing communication techniques for collecting massive distributed sensory data to a central server for completing the downstream tasks. The integration of crowdsensing and communication from a systematic view remains an unexplored area.
Figure 5: Illustration of three ISAC paradigms.
IV-B Integrated Wireless Sensing and Communication
IV-B1 General Integrated Wireless Sensing and Communication
With the explosion of mobile traffic, the available frequency bands for communication are crowded, which drives an urgent demand for more spectrum resources. One promising solution is to reuse the wide radar spectrum [210], giving rise to an emerging research topic i.e., integrated wireless sensing and communication [29]. On the other hand, many radar sensors are deployed for significant applications like air traffic control, geophysical monitoring, surveillance for defense, etc. The rising interference from wireless communication systems should be well managed and suppressed. To tackle these issues caused by communication and radar spectrum sharing, two frameworks are widely investigated, namely, RCC and DFRC [209]. The former aims to develop efficient interference-suppression techniques so that the two systems can coexist without interference. The latter focuses on designing systems that can simultaneously realize wireless communication and remote sensing. In the sequel, the techniques for empowering these two frameworks are elaborated.
The main issue in an RCC system is to avoid interference between radar and communication. Recent research starts from designing opportunistic spectrum access and interference management or cancellation schemes based on interference channel state information (ICSI) estimation. Similar to cognitive radio, radar is regarded as the primary user and the communication system plays the role of the secondary user. To be specific, a communication transceiver intelligently detects whether a communication channel is in use or not and moves transmission into vacant channels. This makes better use of spectrum resources and enhances the frequency efficiency. To avoid the possible performance degradation of radar caused by co-channel interference, the communication system should control its transmit power [211]. Although it is simple to implement opportunistic spectrum access schemes with low cost, there is no real spectrum sharing between radar and communication. Besides, such schemes cannot facilitate the coexistence of communication and MIMO radar, since it’s difficult for the former to identify the sidelobes with radar’s omnidirectional waveforms. Therefore, the most recent work focuses on investigating frequency-sharing schemes under the target of interference cancellation based on ICSI estimation.
The key to support spectrum sharing between radar and communication is ICSI estimation. Conventionally, ICSI is estimated by exploiting pilot signals sent to the radar by the communication system [212]. An alternative method is to set a dedicated coordinator connected to both systems via wireless or backhaul links for ICSI estimation and transmit precoding design [213]. These methods, however, cause additional computational and signaling costs. To alleviate the estimation overhead, the radar probing waveform can be utilized as the pilot signal [214]. Based on the knowledge of ICSI, different techniques are proposed for supporting the spectrum sharing between radar and communication systems, as discussed below.
  • Null-space Projection (NSP) Radar Precoders: In this scheme, the radar precoder is designed to transmit waveforms only in the null space of the interference channel matrix, which is obtained by applying singular value decomposition (SVD) [215]. As a result, the interference from radar to communication is strictly avoided. However, this scheme limits the performance of radar, especially for MIMO radar where the null space of the interference channel is small. To address this issue, some other schemes enhance the radar performance by allowing non-zero interference in the dimensions with small singular values at the communication receiver (see, e.g., [216]).
  • Receiver Design: Most works on this topic focus on accurately demodulating the communication data by suppressing the radar interference. By exploiting the sparse property of the radar interference and the communication error caused by channel fading and noise, compressive sensing is utilized to recover the radar interference and the communication signal simultaneously [217]. For radar using pulsed waveforms, its interference either has high amplitudes in short durations or is zero otherwise, which keeps unchanged for a relatively long time scale due to the slow variation of radar settings. Thus, the interference amplitude can be accurately estimated with low cost. Benefiting from this fact and by assuming the phase shift of radar interference in a uniform distribution in the range of $[0,2\pi]$, the optimal demodulating scheme is first investigated under given modulation constellations, followed by the design of self-adaptive constellations that optimize communication capacity and the symbol error rate [218].
  • Joint Design: In this case, the radar and communication system parameters such as transmit power and beamforming are jointly adjusted. The pioneering work considers the coexistence of a matrix completion MIMO (MC-MIMO) radar and a point-to-point MIMO communication system [219]. It randomly sub-samples the radar receive antenna array to enlarge the null space, which provides more degrees of freedom for the communication system without causing interference to the radar system. Particularly, the communication signal’s covariance matrix and the radar’s antenna down-sampling matrix are jointly optimized. To address the case with signal-dependent clutter, the authors in [213] optimize the radar signal-to-interference-plus-noise ratio (SINR) while guaranteeing communication quality. Furthermore, the joint optimization scheme is proposed in [220] under the assumption of imperfect ICSI.
In a DFRC system, radar and communication can be simultaneously performed and jointly designed via sharing hardware platforms and spectrum resources. However, different from communication systems, radar systems have uncooperative targets and the useful information is in the echo signals [82]. This leads to different design principles. To unify the design of radar and communication, researchers have made efforts to investigate the performance metrics of radar from an information-theoretic view. In [221], the capacity of a round-trip radar channel is defined as the number of distinguishable targets. Furthermore, the authors in [222] point out that the maximum achievable information of detected targets of a radar system is the mutual information of the transmit signal and the echo waveform. Thereby, the target of echo signal processing is to reduce the uncertainty of detected targets for enhancing the detection precision. The remaining part of the uncertainty is lower-bounded by the Cramér-Rao bound (CRB) [223]. On this basis, the performance tradeoff between communication and radar is investigated [222, 224, 225].
Another main research topic of DFRC is to design dual-functional waveforms useful for both radar and communication. The pioneering work utilized pulsed interval modulation to modulate communication bits on the radar pulsed waveforms [83]. On top of this method, many work designed dual-functional waveforms for both communication and radar via different modulation schemes (see, e.g., [226, 227, 228]). Whereas, a better approach is to exploit the existing communication signals for completing sensing tasks. Particularly, the widely used OFDM signals in modern communication systems are regarded as a promising candidate [229]. Fast FFT and IFFT are utilized for estimating the DFS and range information [230] from the OFDM waveforms. What’s more, researchers further target designing dual-functional waveforms in MIMO systems via beamforming. One basic approach is to use the main lobe of the MIMO signal for target detection and transmit information via the sidelobe [231, 232]. However, in this approach, communication symbols are embedded into the radar pulses, and thus the communication data rate depends on the pulse repetition frequency (PRF) of the radar, which is much lower than the conventional communication systems. Besides, communication is not available when the sidelobe has no line-of-sight link to the receiver. To address these shortages, better schemes directly use the communication waveforms for target detection, which optimize the transmit beamforming for balancing the performance tradeoff between radar and communication (see, e.g., [233, 234]). Moreover, a series of tecnologies are exploited to improve the performance of ISAC, such as the intelligent reflecting surface (IRS) [235] and wireless power transfer [236, 237, 238, 239].
IV-B2 Wi-Fi Sensing
As its name suggests, Wi-Fi sensing, as a special case of integrated wireless sensing and communication, exploits Wi-Fi signals for detection, recognition, and estimation [36]. Compared with traditional radar sensing, it has a lower cost since the Wi-Fi signals and equipments are ubiquitous and no extra hardware is needed [37]. Wi-Fi sensing can be widely applied in a series of scenarios such as fall detection [240], indoor tracking [241], and so on. Particularly, OFDM signals are used for sensing, since OFDM modulation is widely adopted for efficient communication in Wi-Fi systems by dividing the frequency-selective broadband channels into multiple orthogonal subcarriers with flat fading in each subcarrier [36]. The essential data for completing Wi-Fi sensing tasks is CSI of OFDM signals, as CSI can describe the target effects on amplitude and phase of the signal and can provide rich information over multiple subcarriers [36]. The collected CSI is first processed before inputted into AI models for finishing the tasks, i.e., detection, recognition, or estimation. In the sequel, the signal processing techniques are described.
The first important signal processing technique is to extract useful features from the raw CSI data. The extracted features include the amplitude and phase of signals, their difference over subsequent time slots to capture the temporal changes as features [242], the statistical features like mean, standard deviation, median, kurtosis (see, e.g., [243]), the signals’ power density in the frequency domain (see, e.g., [244]), or the time-frequency components generated by wavelet transform (see, e.g., [245]). Then, it’s significant to cancel the interference and noise in the received Wi-Fi signals before being used for completing the sensing task, since interference from other sources and reflected by the environmental obstacles is very common in Wi-Fi systems [36]. To cancel the interference, many typical filters are used, including moving average over several time slots (see, e.g., [246]), Hampel filter [247], and FFT frequency filter [248]. Subsequently, the collected CSI sample often has redundant information since signals over adjacent subcarriers have similar contents [249]. To exclude the redundant information for reducing computation overhead of downstream tasks (e.g., detection and recognition), dimensionality reduction is performed, like selecting subcarriers with large temporal variance (see, e.g., [249]), deleting the subcarriers with negative correlation to others (see, e.g., [250]), using PCA to select the principal eigenspace of all subcarriers (see, e.g., [251]), etc.
In addition to the common signal processing techniques, i.e., feature extraction, interference cancellation, and dimensionality reduction, mentioned above, there are many application-specific pre-processing approaches. Specifically, the method of detrending targets mitigates the influence of fluctuating trends caused by environmental change via using approaches like least square regression [252]. Interpolation is applied when there is packet loss in the collected CSI [253]. Segmentation is exploited to decide how many CSI samples should be collected over time for the following decision-making using e.g., a window with a specific start and stop sentinel movement [254].
IV-C Integrated Multi-Modal Sensing and Communication
Although integrated wireless sensing and communication have achieved significant development, such techniques cannot fully utilize the massive non-RF sensors and the corresponding sensory data [38]. What’s more, in future wireless networks, the electromagnetic environment may change rapidly, leading to severe frequency selectivity and time selectivity in wireless channels. The performance gain from integrated wireless sensing and communication, however, comes from the electromagnetic wave propagation domain, which cannot provide either complete or fine-grained characterization of the rapidly-changing environment. This leads to little performance enhancement potential [38]. On the contrary, multi-modal sensing systems obtain information of the environment through multiple domains, including radio wave, visible light or infrared light reflections. With the widely deployed multi-modal sensors, the parameters of the physical environment can be acquired more precisely and efficiently, thereby enhancing communication efficiency. On the other hand, with fast and reliable communication networks, multi-modal sensory data can be shared, exchanged, and fused among different devices, which can further significantly enhance the sensing quality and the performance of downstream tasks such as AI model training and inference. To this end, the research paradigm of integrated multi-modal sensing and communication is proposed [38]. In the sequel, the challenges and techniques are first elaborated. Then, a promising research topic of integrated multi-modal sensing and communication called haptic communications, is introduced [85], which is regarded as a key technology in 6G.
IV-C1 Challenges and Techniques
The efficient implementation of integrated multi-modal sensing and communication faces two main kinds of challenges. On one hand, the sensory data obtained from non-RF sensors are usually in an abstract format. This makes it difficult to be integrated with communication, thereby resulting in a lack of processing methods with interpretability and expansibility. On the other hand, there is currently a lack of datasets that contain aligned multi-modal sensory data and wireless channel information. As a result, the mapping relation between multi-modal sensing and communication remains an unexplored issue.
The techniques of integrated multi-modal sensing and communication can be grouped into three types. The first type targets constructing datasets for building up the relation between multi-modal sensory data and communication environments. The second type aims at exploiting the multi-modal sensory data and machine learning algorithms for CSI estimation. The last type enhances the performance of multi-modal-sensing-based intelligent tasks.
  • Dataset Construction: Since there are various kinds of multi-modal sensors, the obtained multi-modal sensory data are in an abstract format, which results in a lack of processing methods with interpretability and expansibility. Besides, the mapping relation between multi-modal sensing and communication is unexplored. To tackle these issues, a good dataset of integrated multi-modal sensing and communication should satisfy the following requirement [38]. To begin with, the sensory data and wireless channel data should be aligned to achieve mutual promotion between each other. Then, the sensory data should be collected from various devices in order to enlarge the sensing ranges, thereby enhancing the scalability and generality of the supported applications. Moreover, the dataset should be obtained from real scenarios to make sure that its supported research and projects have practical and high-generation ability. To this end, a vision-aided wireless communication dataset is constructed using 3D modeling and ray-tracing simulators [255], where RGB images and depth maps are created from the visual instances, and the wireless features are obtained with sinusoid waveform generator. By exploiting the dataset in [255], a machine-learning-based framework is proposed in [256] for enabling proaction in high-frequency wireless networks. The authors in [257] develop a simulation dataset called M3SC in dynamic vehicular communication networks, where the multi-modal sensory data and wireless channel data are aligned and massive MIMO and millimeter wave (mmWave) communications are considered. For better characterizing the practical imperfections, a large-scale integrated multi-modal sensing and communication dataset is proposed in [258], where the data are collected from real-world scenarios with mmWave receivers, RGB cameras, 3D LiDARs, RF radars and GPS receivers.
  • CSI Estimation for Communication Enhancement: Multi-modal sensory data include information on key elements in the environments such as the position of scatters, that are highly relevant to CSI. Therefore, using multi-modal sensory data for CSI estimation has attracted much focus. However, the mapping relation between multi-modal sensory data and communication characteristics is difficult to obtain. Many works manage to build up such a relation via using deep neural networks (e.g., [84, 259, 260]). Another research direction targets enhancing the communication performance in MIMO systems, where beam should be aligned to increase the received SNR. Multi-modal sensory data has been useful for predictive beamforming, where the position of a target is predicted with the assistance of multi-modal sensors so as to perform the corresponding beamforming in advance (e.g., [261, 262]).
  • Multi-Modal Sensing Ability Enhancement: In multi-agent multi-modal sensing systems, communication plays an important role in data fusion among different devices for e.g., obtaining beyond-vision global and fine-grained sensory data. There are three cross-device multi-modal data fusion schemes. The first is raw data fusion, where raw sensory data is shared among different devices, leading to the largest information redundancy and communication overhead [38, 263, 264, 265]. To overcome the communication bottleneck in raw data fusion, the feature fusion scheme is proposed, where low dimensional feature vectors are extracted and shared at each device [266, 267, 268, 269]. The last scheme is semantic fusion, which requires the least communication bandwidth. In this scheme, the semantic results of different devices are shared [270, 271].
IV-C2 Haptic Communications
Haptic communications refer to exchanging haptic information in networked teleoperation systems for supporting remote immersive applications, including tele-diagnosis, tele-surgery, remote teaching, virtual/augmented reality, and so on [85]. Haptics include multiple modalities of sensory data from both kinesthetic perception (including forces, torques, position, velocity, etc.) and tactile perception (including surface texture, friction, etc.) [85]. Although traditional remote interactions such as voice or video can well support many remote applications like remote voice and video conferencing, the achievement of remote immersion demands additional haptic information [85]. Last but not least, the requirements of high packet rate, low packet loss, and variable delay lead to hard management and synchronization of the data streams [272].
The challenges of haptic communications lie in three aspects. The first is the huge communication load caused by the haptic sensors. To guarantee the stability and user experience of teleoperation systems, many types and quantities of haptic sensors should be deployed and they obtain and transmit data at a high rate, e.g., exceeding 1 kHz [273], which brings huge communication overhead. Besides, remote immersive applications, especially critical missions such as telesurgery, are very sensitive to data loss and latency [274]. In summary, haptic communications demand ultra-high reliability and low latency [275].
Existing techniques focus on improving the communication efficiency for transmitting the haptic sensory data, generally together with the corresponding video and audio data. The correlation among the signals of different modalities can be utilized for data compression thus alleviating communication overhead [276]. A cross-modal communication architecture is proposed in [277], where the heterogeneous modal requirements are satisfied and the cross-modal correlation is utilized for data reconstruction. Besides, data in different modalities has different transmission rate demands [278]. For example, the transmission of audio/video data calls for high throughput, whereas that of haptic data demands ultra-high reliability and low latency [275]. Based on this property, several modality-aware communication techniques are designed via e.g., setting priority for haptic data [279] and adaptive resource allocation among different modalities [280].
IV-D Integrated Mobile Crowdsensing and Communication
Typical mobile crowdsensing services are implemented on central servers or clouds. However, the collection of a huge amount of sensory data from massive mobile devices causes a large communication overhead. Besides, most mobile devices communicate via hostile wireless links with channel fading, noise, and limited network resources. Furthermore, many crowdsensing services demand real-time decisions and fast responses, resulting in more severe communication bottlenecks. To maximize the overall crowdsensing system utility calls for communication techniques to overcome the above drawbacks. A direct approach adopts the existing communication schemes in telecommunication networks. They are divided into the edge-network-level techniques and the core-network-level techniques [33]. The former collects sensory data in local wireless edge networks via utilizing existing protocols, including ZigBee, Bluetooth, WiFi, and cellular communications. The latter targets remote data transmission e.g., from mobile devices to clouds, by using a high-speed fiber-optic architecture to provide connections among different wireless edge networks and clouds. Although directly adopting existing communication schemes has low design complexity and is cost-effective, the separated design between communication and crowdsensing degrades the resource utility and cannot fully unleash the performance potential [281]. To this end, a joint design that integrates crowdsensing and communication from a systematic view is desirable. There is currently little work focusing on this area, leaving many research opportunities.
V Signal Design for ISCC
V-A Overview
The above partial-integration techniques, i.e., ISC, ICC, and ISAC, fall short in three aspects as mentioned in Section I-B, i.e., low resource utilization caused by partial integration, mismatch between goals of the overall task and each module caused by separate designs, and ignoring the tight coupling of different modules. To overcome these shortages for improving system performance, ISCC designs are desirable.
To support ISCC, the signal designs have been investigated including the beamforming and waveform designs. Specifically, the waveforms need to be designed to balance the radar sensing performance evaluated by CRB, the communication performance evaluated by SNR, and the computation performance evaluated by AirComp mean-squared-error (MSE) between the aggregated data and the ground truth. The beamforming designs are further required to equalize the channels among mobile devices and suppress noise to improve the ISCC performance.
In the physical layer, ISCC can be categorized into three types as shown in Fig. 6. The first type designs independent signals dedicated for sensing, communication, and computation functionalities, which is known as the single-functional signal design. However, the dedicated signals cause severe interference among different functionalities. To reduce the interference, the second type designs one signal for sensing and another for AirComp, which is defined as the dual-functional signal design. As a further step, the third type namely the triple-functional signal design realizes sensing, communication, and computation in one signal, which trades increased design complexity for interference cancellation. The technical details are elaborated as follows.
Figure 6: ISCC signal designs.
V-B Single-Functional Signal Design
In [86], the independent communication signals and computation signals are transmitted simultaneously from multiple multi-antenna mobile devices (MDs), while the data is assumed to be sensed by the MDs. The communication signals are received by the server to recover the data with the existence of interference and channel noise, where the performance is evaluated by the signal-to-interference-plus-noise ratio (SINR). As for computation, The signals are superposed during the transmission and aggregated at the server, where the performance is evaluated by the MSE between the aggregated data and the ground truth. To improve the communication throughput and reduce the computation error, the transmit and receive beamforming designs for communication and computation are jointly designed. However, the neglect of the sensing process makes it less possible to be implemented in practice.
By taking the sensing process into consideration, the dedicated signals for sensing, communication, and computation are transmitted simultaneously from multiple single-antenna MDs in [282]. The sensing signals are reflected by the targets and then received by the server, based on which the information of targets is estimated. The performance of sensing is evaluated by the MSE between the estimated target information and the ground truth. The computation signals carrying local learning results at each sensor are aggregated at the server to update the global learning model, where the performance is evaluated by the MSE between the aggregated local model parameters and the ground truth. To reduce the interference between different functionalities, dedicated beamforming designs for sensing, communication, and computation have been investigated. However, the interference among different signals still exists. To reduce the interference, a series of works explore the possibility to integrate the functionalities of sensing, communication, and computation in one common signal.
V-C Dual-Functional Signal Design
In the dual-functional signal design, the antenna array on each MD is divided to produce two distinct signals, one for sensing purposes and the other for AirComp [45]. The information of the sensing target is extracted from the echo signal received at the MD. At the same time, the gathered data is computed through the analog addition of signal waves during the data transmission from the MDs to the server. The MSE of the estimated target response matrix is used for evaluating the sensing performance, while the MSE of the computation results is used for evaluating the AirComp performance. To improve the AirComp accuracy while guaranteeing the sensing performance, the sensing transmit beamformers, data transmit beamformers, and data aggregation beamformers are jointly designed. The impact of AirComp signal interference on radar sensing can be mitigated using statistical methods, while the interference of sensing signals on AirComp will result in deteriorated performance, which needs to be further dealt with.
V-D Triple-Functional Signal Design
As a further step of integration, ISAC and AirComp are combined in a triple-functional signal design, which is known as the integrated sensing, communication, and computation over-the-air (Air-ISCC) [46]. At each MD, the entire antenna array produces a single signal to realize sensing, communication, and computation functions. The signal sent by each MD serves as both the radar probing pulse and the data carrier, necessitating the design of a transmission beamformer on each MD. The target of radar sensing can be determined from the statistical data of the reflected signal using maximum likelihood estimation (MLE). As for AirComp, the data aggregation beamformer is deployed at the server for equalizing the received signals. Compared with the single and dual-functional signal designs, the triple-functional signal design results in a more complex optimization problem for minimizing the AirComp error while guaranteeing the sensing performance. The desired design is achieved using a solving approach that employs semi-definite relaxation.
To improve the performances of radar sensing and AirComp, a series of beampattern designs are further investigated in [283]. Both the omnidirectional and directional radar beampatterns are designed, where the AirComp accuracy of the directional beampattern design is not as good as that of the omnidirectional one. Additionally, the balance between the optimal radar sensing beampattern and AirComp accuracy is explored. Accounting for the total power and per-antenna power budget constraints respectively, the corresponding beampattern designs for balancing the trade-off are proposed in [284].
As a promising technique for improving spectrum efficiency, Air-ISCC can be applied in a series of scenarios including target localization, vehicular networks, and federated learning [285]. In federated learning, the Air-ISCC signals are used to simultaneously sense the information of the target for local model training and deliver the local training results to the server via AirComp for global model update [286]. Moreover, an IRS is further deployed to assist the AirComp [287].
VI Network Resource Management for ISCC
VI-A Overview
Apart from signal designs for ISCC, another main research stream of ISCC aims to enhance the network resource utility via resources (e.g., time, bandwidth, and energy) sharing. As shown in Fig. 7, there are two ISCC resource allocation paradigms. One paradigm is joint RRM for task coexistence, which allows the coexistence of sensing, communication, and computation tasks, leading to loose coupling among the three modules as shown in Fig. 7(a). In this paradigm, each task has its own goal. Dynamic network resource allocation among different tasks can be conducted to achieve multiple goals (e.g., communication capacity, computation accuracy, and sensing SNR) simultaneously. The other paradigm is task-oriented ISCC resource allocation for complicated tasks that involve sensing for information acquisition, computation for information processing, and communication for information transmission. In these tasks, the three modules are tightly coupled as shown in Fig. 7(b). The reasons are as follows. On one hand, the overall task performance depends on the qualities of the three modules, e.g., the sensing and communication SNR, the computation latency, and the quantization error. However, they should compete for resources to enhance their own qualities. On the other hand, there is tight dependency among them in general. For example, the size of the obtained sensory data decides the computation load, while the size of on-device computational results further determines the communication load. In the sequel, the above two paradigms are elaborated.
Figure 7: Two ISCC resource management paradigms.
VI-B Joint RRM for Task Coexistence
In this paradigm, sensing, communication, and computation tasks coexist and network resources are jointly and adaptively allocated among them to achieve multiple goals at the same time. In general, sensing and communication tasks are executed in physical layers, while computation tasks are realized in upper layers by e.g., MEC. In this paradigm, each type of task adopts its traditional performance metric. As an example, in [288], the sensing performance is measured by the mutual information (MI), the communication task aims to achieve a higher transmission rate, and the computation performance is evaluated by the processing delay. In the following, three typical example cases of this paradigm are discussed, i.e., the coexistence of sensing, communication, and computation tasks, the coexistence of sensing and MEC tasks, and the coexistence of ISC and MEC tasks.
VI-B1 Coexistence of Sensing, Communication, and Computation Tasks
In this case, sensing, communication, and computation tasks coexist and compete for network resources to achieve their own targets. Particularly, in [288], ISAC serves as a supporting technique to allow hardware and spectrum sharing between the communication and the sensing tasks. A Cobb-Douglas utility function that combines three metrics, i.e., sensing precision, communication capacity, and computation capability, is adopted and maximized by the joint design of device association and sub-channel assignment. Besides, the authors in [289] investigate a time-division multiplexing ISCC system, where FMCW is used as the sensing waveform and orthogonal time slots are assigned to the sensing, uplink/downlink communication, and computation processes. Targeting communication capacity maximization and computation offloading latency minimization while guaranteeing the sensing distortion is under a limited threshold, a joint resource allocation, time partitioning scheme, computation task processing mode selection, and target sensing location selection, are proposed. In [290], the IRS is applied to assist the ISAC and computation task offloading, where both partial offloading and binary offloading are considered. A performance metric called weighted throughput capacity, which comprises the radar sensing capacity, communication throughput, and computation throughput, is proposed and maximized via joint beamforming design and time allocation. Furthermore, in [28], the ISCC framework is designed for intelligent machine-type communication (IMTC) networks to achieve highly reliable communication with low latency, highly accurate sensing, and fast data processing. In IMTC, intelligent machines (IMs) are expected to execute complex data processing tasks in rapidly varying environments, which demands IMs to conduct real-time environmental sensing and cooperation with each other [291]. To this end, a closed-loop is formulated where the sensed data is exploited as a priori information to improve the communication performance, the communication techniques are designed to support the intelligent computation, and the computation results in turn facilitate the ISAC process.
VI-B2 Coexistence of Sensing and MEC Tasks
In this case, the sensing and the MEC tasks compete for radio resources to achieve their respective goals. ISAC technique is used to enable hardware and radio resource sharing between sensing and communication. In [88], a UAV-supported ISCC framework is proposed, where a treble-functional UAV senses a target by using the unified beam waves, processes sensory data locally, and offloads its computational tasks to ground APs. By characterizing the Pareto boundary between the sensing beampattern gain and the computation capacity, a fundamental trade-off is revealed. Specifically, by applying the semidefinite programming method and the concave-convex procedure technique, a joint optimization algorithm is proposed and depicts the performance region, as well as enhancing the ability of sensing, computing, and communication of UAVs. Besides, an architecture that integrates communication, radar sensing, and MEC, is developed in [292], where user terminals conduct computation offloading and radar sensing over the same radio resource blocks by using DFRC techniques and MIMO arrays. To jointly optimize the performance of MIMO radar beampattern design, computation offloading energy consumption, the individual transmit precoding for radar, and resource allocation of communication and computation, a multi-objective optimization problem is formulated. Moreover, the authors in [293] develop a non-orthogonal multiple access (NOMA)-aided joint communication, sensing, and multi-tier computing (JCSMC) framework, where the BS can not only sense target areas and offer MEC services to the covered mobile devices simultaneously but also offload some computation tasks to cloud servers to improve computation efficiency. The potential benefits of utilizing NOMA in this framework are investigated. Both the transmit beamformer and computing resource allocation are jointly optimized to maximize the computation rate in two computation offloading modes, i.e., partial and binary offloading modes.
VI-B3 Coexistence of ISC and MEC Tasks
In this type, an ISC task competes with an MEC task for network resources to enhance respective computational performance. Particularly, in [87], computation resource competition between an AI model inference task and an MEC offloading task in an edge server is investigated where sensing serves as the supporting technique for completing the AI model inference task and communication serves as the supporting technique for the MEC task. A novel action detection module is first proposed for alleviating the network resource competition by detecting the status of the sensing target. If the sensing target is static, no further computation is required for the AI model inference task and hence reduces the communication overhead at the edge server. Then, the inference accuracy is maximized while guaranteeing the quality-of-service (QoS) requirements of the MEC task via designing an optimal strategy of joint resource allocation and static status threshold selection.
VI-C Task-oriented ISCC Resource Allocation
In this paradigm, network resources are allocated among the sensing, communication, and computation modules of a complex task to achieve its customized goal. The fulfillment of complex tasks requires the cooperation of sensing for information acquisition, communication for information transmission, and computation for information processing, such as auto-driving tasks, federated learning tasks, edge inference tasks, etc. The qualities of the three modules decide the performance of these tasks and they compete for the limited radio resources. Moreover, there is a lack of a uniform design objective for these tasks. For example, instead of targeting throughput maximization in conventional communication tasks, a federated learning task concerns improving learning accuracy while reducing training latency. Therefore, various kinds of task-oriented ISCC resource allocation schemes are developed, as discussed below.
VI-C1 ISCC for Vehicular Networks
In this case, task-oriented ISCC resource allocation schemes are proposed to enhance the system performance, e.g., achieve low latency for auto-driving [89] and address road traffic congestion [294]. Specifically, in [89], the message and resource allocation are investigated for cooperative perception in fog-based vehicular networks. The satisfaction of cooperative perception (SCP) is firstly characterized based on the spatial-temporal value and latency performance. To maximize the sum SCP, the sensing block message, communication resource block, and computation resource are jointly allocated under the maximum latency and sojourn time constraints of vehicles. Besides, in [294], the ISCC technique is regarded as an appealing control tool for road traffic congestion. The evolution of road traffic congestion control is divided into two sequential stages, i.e., the sensing and the communication stages. In the former stage, congestion control strategies experience rapid growth due to the accessibility of traffic data. In the latter stage, with advances in vehicular networks, big data and AI, the improvement of the ability to collect finer data is significant. This paper also introduces how existing congestion detection techniques evolve with the continuous development of sensing, communication, and computation capability in the networks.
VI-C2 ISCC for Machine Communications
In on-purpose machine communication (OMC), the interactive information flow is the machine semantics, which refers to the consensus information reached by the intelligent machines (IMs) after feature extraction and transformation from the raw data [90]. A cross-layer ISCC framework is designed for OMC in [295], where the ISCC is deployed at the cloud, edge, and terminals, and the IMs can transmit directional beams to the dedicated devices according to their communication purposes
VI-C3 ISCC for Mobile Crowdsensing
In [75], mobile crowdsensing is modeled as a typical sensing, computation, and communication process. Due to hostile wireless channels, limited spectrum, energy, and computation resources, the transmission or processing of sensing data may be unsuccessful in wireless edge networks. Thus, an ISCC framework for multi-dimensional resource-constrained crowdsensing systems is proposed to optimize the overall performance by jointly selecting users, allocating bandwidth, and designing the strategies.
VI-C4 ISCC for Sensor Networks
In [91], a semantic model and a hybrid perceptive frame approach are proposed for the control of ISCC-based reconfigurable mobile sensor networks. First, the mobile sensor network is modeled as an ISCC-based dynamic system. Then, the relation between different sensors and the corresponding multi-modal sensory information is characterized by a proposed semantic model. On this basis, by describing the hybrid nature of cooperative sensor networks and the integration of sensing, communication, and cooperation as perceptive references and semantic representations, the mobile sensor network can autonomously reconfigure itself adapting to the varying operating conditions.
VI-C5 ISCC for Federated Learning Tasks
In [48], the online federated-learning-guided ISCC schemes are proposed. The sensing, on-device computation, and communication operate in a time-division manner. The FMCW radar signals are transmitted by the MDs in the sensing mode. The radar echo signals are received by the MDs and locally processed to extract the target information in the computation mode. The locally processed data is then delivered from MDs to the server for global model updating in the communication mode. Aim at maximizing the convergence speed of model training, the sensing, communication, and computation modules are carefully designed. However, the competition for time resources will limit the performances of these modules and thus the learning performance.
VI-C6 ISCC for Edge AI Inference Tasks
For real-time edge inference tasks, a task-oriented AirComp scheme is first proposed in [26]. In this work, the AI model is split into two parts, the front-end part deployed at devices is used for feature extraction from sensory data. The backend part with intensive computation deployed at an edge server completes the inference task by receiving all local feature vectors from all devices. Specifically, multiple devices sense the same wide view of the source target and obtain a local distortion-corrupted version of the ground-true sensory data. Then, each device extracts a low-dimensional local feature vector from its local sensory data using a principal component analysis based linear extractor. An edge server aggregates all local feature vectors via AirComp to obtain a denoised global feature vector, which is input into the remaining AI model for finishing the downstream inference task. This work directly adopts the inference accuracy as the design objective rather than the conventional metric of MMSE, as the latter ignores the feature elements with the same distortion but heterogeneous contributions to the inference accuracy. Besides, the communication efficiency is enhanced using AirComp. The above work is further extended to the case with sensing noise suppression [296] and the decentralized case [297]. Besides, to deal with the scenario where different devices sense disjoint narrow views of the source target, a task-oriented sensing, feature quantization, and feature transmission technique is proposed in [47], where the local feature vectors of different devices are cascaded at the server to form a global one. Furthermore, a device scheduling based ISCC approach is proposed in [298] for enhancing the inference accuracy while preserving the data privacy.
VII ISCC in Future Advanced Networks
In this section, the potential ISCC techniques for deploying digital twin platforms at the network edge and the potential ISCC applications in two future advanced networks, i.e., the computing power networks and the SAGINs, are discussed.
Figure 8: An ISCC-based digital twin network.
VII-A ISCC for Digital Twin Enabled Wireless Networks
A digital twin system refers to constructing virtual or digital representations of the physical objects and systems via the techniques of high-quality distributed sensing, precise 3D map construction, real-time tracing, deep machine learning, and so on [299]. It targets generating simulation data that cannot be distinguished from its real-world counterparts, conducts data analysis, and makes decisions based on the generated data [300]. Digital twins have been widely used in many areas e.g., intelligent manufacturing, climate change, healthcare, and economic crises [300]. Particularly, since digital twins have a powerful ability to simulate complex wireless environments, it has attracted more and more attention to deploy digital twins at the network edge for addressing many issues including network complexity control [301], network performance enhancement [302], disruptions prediction and prevention [303], network topology visualization [304], and supporting MEC and industrial IoT networks [299].
However, establishing a digital twin network is challenging. It desires massive sensory data of multiple modalities for acquiring ground-true information of real wireless networks, high communication throughput for transmitting sensory data or the corresponding extracted information, and powerful computation capability for data generation, analytics, and decision-making. To tackle these challenges, an ISCC-based digital twin network is proposed as shown in Fig. 8, where many sensors of different types are deployed for perceiving a real wireless network to obtain ground-true information, and the digital twin system is deployed at an edge server for executing all complex and computation-intensive algorithms to construct the virtual representation. Based on such a digital twin network, two ISCC techniques are discussed. One is a hybrid learning platform for training AI models that are used for constructing digital twin systems. The other is a real-time ISCC technique to leverage the established digital twin for making accurate and fast decisions.
VII-A1 Hybrid AI Training Based Digital Twin Network Construction
One significant step for establishing the digital twin network in Fig.8 is training many AI models for accomplishing many intelligent tasks, by using a large amount of multi-modal sensory data from many distributed sensors. However, different sensors have heterogeneous resources like energy and different capabilities in terms of sensing, communication, and computation. Besides, different sensors have different data privacy preservation requirements. To deal with this heterogeneous digital twin network, a hybrid AI training platform that integrates federated learning, federated split learning, and centralized learning, is proposed to meet various requirements of different sensors, as shown in Fig. 9. For example, for the sensors with critical privacy concerns and high computation and communication capabilities, federated learning is adopted. For the sensors with high privacy preservation requirements and low computation and communication abilities, federated split learning is performed. For the sensors without privacy concerns and with limited communication and computation resources, they can directly send the raw sensory data to the edge server for centralized learning.
Figure 9: An ISCC-based hybrid AI model training platform for establishing digital twin.
In each training iteration, the three training schemes are conducted in parallel. In federated learning, the global model is shared to the corresponding sensors for local training based on the obtained online and offline local data samples. In federated split learning, the global model is partitioned into a front-end part and a back-end part. The front-end part is shared to all involved devices for executing the forward pass using the online and offline sensory data and all local front-end parts are then transmitted back to the server for completing the downstream training. The centralized learning is conducted at the edge server by continuously collecting the real-time raw sensory data from the corresponding sensors. At the end of the training iteration, the updated models of the three training schemes are further aggregated to update the global model.
The design object of such hybrid AI model training is to minimize the long-term resource consumption (e.g., energy consumption) of all devices to extend their lifetimes under the constraints of high learning accuracy, low training latency till convergence, and the various privacy requirements of different sensors. For simultaneously training multiple AI models in the hybrid platform, resources (e.g., time, energy, bandwidth) allocation needs to be performed among different tasks via joint design of sensing, communication, and computation.
VII-A2 ISCC Based Digital Twin Network Application
One key function of deploying digital twins at the network edge is making accurate and fast decisions for many network tasks such as resource allocation, multiple access, and computation offloading. To achieve this goal, it’s desirable to collect the real-time sensory data from the sensors. Furthermore, the obtained sensory data should be sufficient and important to accomplish the specific task, as well as having high quality which means it cannot be polluted during the sensing, transmission, and processing. To this end, ISCC techniques should be designed, that should jointly consider many factors. One is the sensor’s sensing capability including sensing range, data quality, sensing form (e.g., in an active way like radar sensors or in a passive way like cameras), the importance level of the obtained data to the task, etc. Other factors include the computation and communication abilities, the channel gains, battery lives, maximum transmit power, and the permitted latency for finishing the task.
For simultaneously completing multiple decision-making tasks, the ISCC design is more complicated since different tasks compete for network ISCC resources. For saving network resources, an opportunistic sensor scheduling scheme can be designed to select the sensors, that can obtain real-time and high-quality data that is important to all tasks.
VII-B Computing Power Networks Supported ISCC
Benefiting from the rapid advancement of cloud computing, MEC, and smart devices, ubiquitous computation resources are deployed in the networks. Traditional network architectures are designed from the perspective of high communication efficiency and cause computing power island effect. To overcome this shortage and to better leverage these distributed computation resources, a new computing network paradigm, called computing power networks, is proposed [13, 14]. It aims at coordinating the ubiquitous computing nodes in networks for providing more flexible computing power scheduling. For adapting to the heterogenous computing capabilities of different clouds, edge servers, and mobile devices, the micro-service architecture utilizes fine-granularity computing blocks to build up different complex tasks, where the computing blocks can be assigned to different devices according to their computing powers [305, 306, 307]. For example, a micro-service based video analytics framework is proposed in [308] for saving the computation overhead, where different applications like criminal tracking, traffic monitoring, and urban management, can share universal basic computing blocks for processing a same video, including frame decode, object detection, etc. Based on the micro-service architecture, two potential ISCC applications are investigated in computing power networks as follows.
Figure 10: Comparison of ISCC applications in monolithic computing and micro-service architectures
VII-B1 Computing Sharing for Various ISCC Tasks
As shown in Fig. 10, an edge-device cooperative framework is deployed for supporting several ISCC tasks, where sensing and some basic sensory data processing for feature extraction are conducted at device and the deep neural networks based complicated computation tasks are performed at edge server. As a result, this framework can enjoy three-fold benefits of privacy preservation via avoiding raw sensory data transmission, low communication due to low-dimensional feature extraction, and computation offloading. In Fig. 10, two computing architectures are compared for completing multiple ISCC tasks. In the conventional monolithic computing architecture as shown in Fig. 10(a), each task has its own individual sensory data processing streams. The intermediate computing results cannot be shared among different tasks. Consequently, it leads to a severe waste of computing resources as well as increasing storage and communication costs. To tackle these issues, the micro-service architecture divides the signal processing of each task into many fine-granularity computing blocks as shown in Fig. 10(b). Different tasks can share the same computing blocks, e.g., the blocks of signal sampling, clutter cancellation, and feature extraction for wireless sensing signal and the blocks of stream processing, frame analysis and content identification for video. This significantly saves the computing and storage overhead as well as reduces the communication load for transmitting the on-device processing results to the edge server.
Figure 11: A directed graph model of computing power networks supported ISCC tasks.
VII-B2 Directed-graph-based Joint Scheduling of Sensing, Communication, and Computation
In computing power networks, the fine-granularity computing blocks for completing to make full utilization of the computing resources in various devices. Thereby, the completion of ISCC tasks requires cooperation of multiple nodes. For enhancing the performance of ISCC tasks while saving the resources in terms of sensing, computation, and communication, a directed graph model can be constructed as shown in Fig. 11. In the graph, there are two kinds of devices or called nodes. One is the sensing nodes with the capability of sensing, on-device computation, and communication. The other is the computing nodes with the ability of communication and computation. A target can be sensed by multiple sensing nodes and same computing blocks are deployed at different computing devices. As a result, the fulfillment of an ISCC task has multiple choices of paths. To be specific, a directed edge from node $a$ to node $b$ means that node $a$ transmits its computing results to node $b$ for downstream processing. In each path, sensing and all required computing blocks are sequentially performed for completing the task. Whether there is a directed edge between two nodes depends on whether a wireless communication link can be constructed between them for reliably transmitting the intermediate computing results based on the factors like transmit power, bandwidth, time, and channel gain. In order to enhance the task performance while saving the sensing, communication, and computation costs, it’s desirable to find the best path in the directed graph.
VII-C Space-Air-Ground Integrated Networks Supported ISCC
By integrating satellite, aerial, and ground communication systems, SAGIN has been a promising solution for providing infrastructure-free environments with communication and intelligent services such as precision agriculture, environmental monitoring, and wildlife monitoring [15]. On one hand, compared with conventional terrestrial networks, SAGIN can provide global coverage especially for remote areas by using UAVs as relays and links between UAVs and satellites for providing backhaul connectivity. On the other hand, satellites are equipped with many remote sensors for detecting and monitoring the characteristics of earth systems from the space in a global perspective. This enables real-time and accurate decision-making for many global mission critical such as disaster assessment. Based on the satellites’ two types of functions mentioned above, two SAGIN-supported ISCC frameworks are introduced below.
Figure 12: Terrestrial sensing based ISCC framework in SAGINs.
VII-C1 Terrestrial Sensing based ISCC framework in SAGINs
Consider a hierarchical SAGIN as shown in Fig. 12, where UAVs relay the sensory data from many terrestrial sensing devices to satellites for completing real-time intelligent tasks. Since computation migration between different satellites is difficult due to the high communication load and weak channel link gains between different satellites, only one satellite is selected for completing one task. Each UAV accesses several sensing devices within a disjoint coverage. The communication links between UAVs and the terrestrial sensing devices in their coverage are strong but the links between UAVs and satellites are weak due to the large path loss caused by the long distance. To tackle this issue, all UAVs collect sensory data from the sensing devices in their own coverages, pre-process the collected data for feature extraction and compression to save communication overhead, and forward the extracted and compressed feature. The satellite conducts a global fusion of the extracted and compressed features from all UAVs for executing the task. Compared with traditional terrestrial networks, SAGINs have the ability to connect the sensing devices in remote areas and can support much larger sensing ranges for more complicated ISCC applications.
As well as dealing with the high coupling among sensing, communication, and computation, there are several novel challenges for implementing the above SAGIN-supported ISCC framework. To begin with, satellites have high mobility and the path loss between UAVs and satellites is large, leading to weak and varying communication channels. As a result, new communication techniques are desirable. Next, a satellite has high coverage, resulting in large distance variations between the satellite and different connected UAVs. Consequently, communication time synchronization turns to be a severe issue. Besides, the issues of UAV trajectory design and satellite selection are also significant. Hence, the exploration of such a framework remains an open issue and calls for further efforts.
Figure 13: Remote sensing based ISCC framework in SAGINs.
VII-C2 Remote Sensing based ISCC Framework in SAGINs
Another main function of satellites is remote sensing by being equipped with different types of sensors. As shown in Fig. 13, by obtaining different modalities of sensory data such as panchromatic images, multispectral images, thermal imaging infrared data, and radar sensory data, and transmitting the processed features to the ground station, many real-time tasks can be accomplished such as precision agriculture and climate change detection. To establish such systems, joint design of multi-modal sensing across multiple satellites, on-satellite processing, and satellite-ground communications are required. As well as addressing the same challenge of weak and fast varying channel links as the terrestrial-sensing-based ISCC framework, there are new design requirements including satellite scheduling for selecting the best sensing views and areas of the task, the multi-modal data alignment across different satellites, and the multi-modal data fusion schemes. For example, the data fusion schemes can be categorized into three types. The first is that satellites directly transmit the raw sensory data to the ground station. This scheme can keep the most sensing information but encounters severe communication bottlenecks for transmitting high-resolution sensory data. The second scheme is called early fusion, where each satellite transmits an extracted low-dimensional feature map to the ground station by saving communication load but sacrificing a part of the information. The last one is late fusion, where on-satellite decisions are made and only the results are fused at the ground station via e.g., a majority vote. In this scheme, only very little information is shared to the ground station but high communication efficiency can be achieved. Besides, since late data fusion shares minimum information to the ground station, it can be conducted when satellites and the ground station belong to multiple institutions.
VIII Conclusion
ISCC, with the ability to improve network resource utility and achieve customized goals of complicated tasks, turns to be a key technology for realizing connected intelligence in future networks. In this paper, we first provide a comprehensive survey in terms of the historical partial integration technologies, i.e., ICC, ISC, and ISAC. Then, a holistic discussion on the shortages of these technologies is conducted to motivate ISCC, including low resource utilization, mismatch between goals of the overall task and each module (i.e., sensing, communication, or computation), and ignoring the tight coupling of different modules. Subsequently, the new challenges of ISCC is discussed, including inference management, high signal design complexity, complicated resource management, and lack of unified design criteria. To tackle these challenges, the techniques of signal design and network resource management for ISCC are surveyed. Finally, the implementation of ISCC in several future advanced networks is investigated. Hopefully, this article will promote the advancement of ISCC in 6G, encouraging future research efforts in this interesting field.